### v1

# Compressed and Smooth Latent Space for Text Diffusion Modeling

Abstract
Autoregressive language models dominate modern text generation, yet their sequential nature introduces fundamental limitations: decoding is slow, and maintaining global coherence remains challenging. Diffusion models offer a promising alternative by enabling parallel generation and flexible control; however, their application to text generation is hindered by the high dimensionality of token-level representations. We introduce Cosmos, a novel approach to text generation that operates entirely in a compressed, smooth latent space tailored specifically for diffusion. This space is learned using an autoencoder trained simultaneously for token-level reconstruction and alignment with frozen activations from a pretrained language encoder, providing robust semantic grounding and enabling effective perturbation-based augmentations. Empirically, we demonstrate that text representations can be compressed by 8×8$\times$8 × while maintaining generation quality comparable to token-level diffusion models. Furthermore, increasing the latent sequence length allows Cosmos to surpass both diffusion-based and autoregressive baselines. We evaluate Cosmos on four diverse generative tasks including story generation, question generation, summarization, and detoxification and compare it with various generative paradigms. Cosmos achieves comparable or superior generation quality while offering more than 2×2$\times$2 × faster inference.
1 Introduction
Autoregressive (AR) language models are a de facto gold standard for text generation (radford2018improving, ; radford2019language, ; brown2020gpt3, ). By factorizing the probability of a sequence into a product of conditional token probabilities, they transform the global generation task into a series of local next-token prediction tasks that can be optimized efficiently via teacher forcing (williams1989learning, ).
The same sequential factorization, however, creates several structural bottlenecks. First, decoding is inherently sequential: the time to produce a sequence grows linearly with its length because each token must await the completion of its predecessor. Second, the process suffers from exposure bias: a single early mistake contaminates the context for all subsequent predictions and cannot be corrected without restarting generation (ranzato2015sequence, ; bengio2015scheduled, ; he2019exposure, ). Third, maximizing local log-likelihood encourages the model to privilege fluency over factual accuracy, resulting in the well-documented phenomenon of hallucination (maynez2020faithfulness, ; ji2023survey, ). Finally, because decisions are made token-by-token, the model lacks an explicit global plan and therefore struggles to maintain long-range logical or narrative coherence (ye2024beyond, ).
In stark contrast, computer vision has been reshaped by diffusion models (ho2020ddpm, ; nichol2021improved, ), particularly by latent diffusion (rombach2022ldm, ; blattmann2023stable, ; andreas2023, ), which first compresses a high-resolution image or video into a compact latent representation and then performs the diffusion process in that space. Operating in a low-dimensional latent manifold reduces computational cost by orders of magnitude and enables breathtaking advances in image and video synthesis. Yet recent works (kouzelis2025eq, ; skorokhodov2025improving, ) also show that diffusion is highly sensitive to the geometry of the latent space: poorly designed latent representations can destabilize training and degrade sample quality, underscoring the need for principled mechanisms to construct diffusable [^1] Diffusable representations refer to latent spaces that permit effective modeling by a diffusion process. representations.
Motivated by these insights, we revisit the foundations of text representation. We hypothesize that the token-level encoding used in contemporary language models is heavily overparameterized for the purpose of sequence generation. Building on the successes of latent diffusion in vision, we pose the following question: How far can we compress textual information into a compact latent space while still matching or even exceeding the generative fidelity of conventional token-level representations?
To answer this question, we develop an autoencoder that maps text into a lower-dimensional space and then train a diffusion model directly in that space. We show empirically that naively minimizing token-reconstruction loss yields a brittle latent geometry that hampers diffusion, whereas introducing robustness- and smoothness-oriented objectives produces a well-behaved manifold conducive to high-quality diffusion synthesis. Our experiments demonstrate that with the right training regime, latent diffusion not only rivals but in several settings surpasses traditional token-level baselines.
These findings challenge the prevailing assumption that token-level autoregression is indispensable for language generation and position latent-space diffusion as a powerful alternative paradigm for future large-scale language models. Our key contributions are as follows:
  • We propose Cosmos — a training recipe for a COmpressed and SMOoth latent Space, that allows to train a diffusion model in more compact latent space while matching the quality of token-level diffusion baselines.
  • We demonstrate that modestly scaling the number of latent vectors enables latent-space diffusion to surpass both token-level diffusion and autoregressive models on an unconditional text generation task.
  • On standard text-generation benchmarks, the proposed latent diffusion achieves up to $\mathbf{2\times}$ faster sampling than conventional token-level diffusion models, while matching or slightly exceeding them in quality and diversity.
2 Related work
Early efforts to bring latent diffusion models into natural language generation treat the diffusion module mainly as a pre-processor that supplies conditioning vectors for an autoregressive decoder. LD4LG (lovelace2023latent, ) trains a diffusion model on compressed hidden states from BART, then feeds the generated latent vectors into the BART decoder to produce text. PLANNER (zhang2023planner, ) adopts a similar two-stage recipe: a fine-tuned BERT encoder produces a 16-token variational latent code, a diffusion model refines that code, and a GPT-2 decoder generates text from this code. While both systems improve controllability, their heavy reliance on powerful autoregressive decoders makes it difficult to isolate and evaluate the intrinsic generative capacity of the latent-space diffusion itself.
A complementary line of work applies diffusion directly to the continuous embeddings of pretrained encoders. TEncDM (shabalin2025tencdm, ) demonstrates that full-length BERT representations can be modeled with Gaussian diffusion and decoded into coherent text without an intermediate autoregressive step. Crucially, however, TEncDM retains the original sequence dimensionality, leaving open a question: How aggressively can such representations be compressed before generation quality collapses?
Our study closes this gap. We devise a training procedure that shrinks BERT-level representations by up to $\bf{8\times}$ while preserving, and in some cases enhancing, their suitability for latent-space diffusion.
3 Preliminary
Diffusion models (ho2020ddpm, ; nichol2021improved, ) learn a data distribution by reversing a progressive noising process. Given a clean sample $z_{0}\sim p_{\text{data}}$, the forward dynamic corrupts the input with Gaussian noise whose magnitude is controlled by a continuous time index $t\in[0,1]$:
$z_{t}\;=\;\sqrt{\alpha_{t}}\,z_{0}\;+\;\sqrt{1-\alpha_{t}}\,\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,I),$ (1)
where the noise schedule $\alpha_{t}$ is monotonically decreasing with $\alpha_{0}=1$ and $\alpha_{1}\approx 0$. A neural denoiser $z_{\theta}(z_{t},t)$ with learnable parameters $\theta$ is trained to invert this corruption by minimising the denoising–score–matching objective
$\mathcal{L}_{\text{DM}}\;=\;\mathbb{E}_{z_{0}\sim p_{\text{data}},\,t\sim\mathcal{U}[0,1],\,\varepsilon\sim\mathcal{N}(0,I)}\Bigl{[}\bigl{\lVert}z_{0}-z_{\theta}(z_{t},t)\bigr{\rVert}_{2}^{2}\Bigr{]}.$ (2)
At inference time the model is applied iteratively from $t=1$ to $t=0$, gradually transforming pure noise into a realistic sample.
3.1 Latent diffusion
Latent Diffusion Models (LDMs) (rombach2022ldm, ; andreas2023, ) improve both compute efficiency and quality by learning the diffusion process in a compact latent space rather than in pixel or token space. An autoencoder with encoder $E$ and decoder $D$ is first trained to achieve high-fidelity reconstruction, $\mathbf{\hat{w}}~{}=~{}D\!\bigl{(}E(\mathbf{w})\bigr{)}~{}\approx~{}\mathbf{w},$ where $\mathbf{z}=E(\mathbf{w})$ denotes the low-dimensional latent. Eqs. (1)–(2) yield a diffusion model trained in latent space whose network can be shallower, and faster than a counterpart operating in the original space, yet still achieves comparable perceptual quality after decoding $D(\mathbf{z})$.
In the remainder of the paper we adopt this framework for textual data: we compress contextualised text representations into the latent space, and train a diffusion model that operates within this space.
4 Methodology
4.1 Overview
Frozen text encoder.
We initialise the pipeline with a pretrained contextual encoder $E_{\text{text}}$ (BERT-base (devlin2019bert, ) by default) that remains frozen during all subsequent training stages. For an input sequence of $L$ tokens $\mathbf{w}=(w_{1},\dots,w_{L})$, the encoder produces a matrix of hidden states $\mathbf{h}=E_{\text{text}}(\mathbf{w})\in~{}\mathbb{R}^{L\times d}$, where $d=768$. Each row supplies a semantically rich representation of its corresponding token tenney-etal-2019-bert . These representations provide a high-fidelity starting point for the compression stage.
Compressor.
To distill the variable-length hidden states into a compact set of latents, we employ a Perceiver Resampler architecture alayrac2022flamingovisuallanguagemodel . It is a 12-layer transformer vaswani2017attention in which every block replaces self-attention with cross-attention. Concretely, let $\mathbf{u}\in\mathbb{R}^{N\times d}$ be a set of learnable vectors initialized randomly and kept at a fixed length $N\!\ll\!L$. In each block, these vectors act as queries ($Q=\mathbf{u}W_{Q}$), while keys and values are formed by projecting a concatenation of $\mathbf{u}$ with the text encoder outputs ($K,V=[\,\mathbf{u};\mathbf{h}\,]W_{K,V}$). Cross-attention therefore allows every vector $u_{i}$ gather information from the entire sequence of hidden states and from other vectors $\mathbf{u}$, gradually refining $\mathbf{u}$ into a semantically organized compressed representation. After the final block we obtain a fixed-length latent matrix $\mathbf{z}\in\mathbb{R}^{N\times d}$, which composes latent space for the diffusion process. For better interpretability of results we compress only along the sequence axis. Altering the embedding dimension $d$ would require architectural changes in the diffusion network itself, conflating encoder effects with diffusion capacity and obscuring the variables we aim to isolate.
Latent normalization.
Before starting the Gaussian diffusion process, we estimate global mean and standard deviation $(\boldsymbol{\mu},\boldsymbol{\sigma})\in\mathbb{R}^{N\times d}$ on a held-out corpus, and normalize each latent feature so that it has zero mean and unit variance, $\mathbf{z}\leftarrow(\mathbf{z}-\boldsymbol{\mu})/\boldsymbol{\sigma}$. This step allows us to run variance-preserving diffusion process.
Latent diffusion model.
We train a Gaussian diffusion model is space formed by $\mathbf{z}$, following Eqs. (1)–(2). Because $N$ is small, the diffusion model runs faster in this space.
Decompressor.
Decompressor mirrors compressor in terms of architecture. It expands the fixed-length latent back to a sequence of $L$ vectors $\hat{\mathbf{h}}\in\mathbb{R}^{L\times d}$, capped at $L_{\max}=512$ by default.
Token predictor.
Finally, a linear projection followed by softmax converts each vector in $\hat{\mathbf{h}}$ into a probability distribution over the vocabulary, yielding the generated text.
In our work, we refer to the combination of the text encoder and compressor as the encoder, while the decompressor together with the token predictor constitutes the decoder.
4.2 Learning a compact text latent space
The previous Section 4.1 outlined our end-to-end pipeline: a frozen contextual encoder supplies token representations, a compressor distils them into a fixed-length latent matrix, and a diffusion model operates solely in this latent space. We now zoom in on the compression and decompression stages.
Figure 1: Overview of our training pipeline. A frozen BERT encoder extracts features, which are augmented before compression. A lightweight compressor–decompressor pair is trained with both token reconstruction (CE) and MSE objectives to produce compact and perturbation-resilient latent representations.
4.2.1 How compact can text latents be?
Figure 2: Token-level reconstruction accuracy on Wikipedia (512 tokens) as a function of $N$.
Figure 2 reports token-level reconstruction accuracy on Wikipedia dataset as a function of the latent sequence length $N$. The compressor outlined in §4.1 is trained using a token-level cross-entropy objective. Remarkably, a lightweight compressor equipped with only 12 transformer blocks is able to encode 512-token sequences into only $N=16$ latents with $100\%$ reconstruction accuracy, achieving a 32×32$\times$32 × compression relative to initial 512 hidden states produced by BERT.
These results paint an encouraging picture: if such a compact latent manifold can be modeled generatively, text generation could proceed in a space whose dimensionality is much smaller than that of the original token embeddings, promising dramatic speed-ups. However, as we show further, high reconstruction accuracy by itself does not imply that the latent manifold is suitable for generative modeling.
4.2.2 Learning robust compact text representations
The reconstruction study (§4.2.1) reveals that aggressive text compression can be loss-free. However, for a robust diffusion generation compressed space should satisfy additional requirements. Empirically, we observe that if the latent manifold lacks smoothness and robustness (Section 5), a Gaussian diffusion model fails to sample latents that yield high-quality texts. To improve diffusability of the latent space, we employ three complementary strategies to the autoencoder training.  Figure 1 provides an overview of the proposed approach.
MSE regularisation on encoder activations.
Alongside the cross-entropy loss between the original tokens and their reconstructions, we add a mean-squared error penalty between the frozen text encoder outputs $\mathbf{h}$ and their reconstructions $\mathbf{\hat{h}}$. This auxiliary objective forces the compressor to preserve the semantics carried by the contextual representations $\mathbf{h}$.
Activation-space perturbations.
In order to teach compressor to extract additional features from $\mathbf{h}$ instead of just preserving its semantics, we apply perturb-and-recover training procedure. Concretely, we sample an augmented view $\mathbf{h}^{\prime}$, pass it through the compressor–decompressor pipeline to obtain a reconstructed representation $\mathbf{\hat{h}}^{\prime}$, and minimise $\operatorname{MSE}(\mathbf{h},\mathbf{\hat{h}}^{\prime})$, forcing both compressor and decompressor to remain invariant to the perturbation. Two perturbations are applied with equal probability within every batch:
  1. (a)
    Random masking: $30\%$ of the vectors of $\mathbf{h}$ are set to zero, and
  2. (b)
    Gaussian noise: after normalising $\mathbf{h}$ with the pre-computed statistics, we inject noise via $\mathbf{h}^{\prime}=\delta\,\mathbf{h}+\sqrt{1-\delta^{2}}\,\varepsilon$ with $\delta=0.7$ and $\varepsilon\sim\mathcal{N}(0,I)$.
These augmentations force the autoencoder to tolerate partial information loss and promote smooth interpolation between adjacent latents (see Section 5.1).
Latent-space augmentation.
We also apply augmentation directly to the latent matrix $\mathbf{z}$. Since the number of latent vectors $N$ can be an order of magnitude smaller than the token count $L$, masking whole latent vectors would annihilate too much information. Instead, during training we randomly zero out a fixed proportion $p$ of the individual features inside each latent vector. This fine-grained sparsification encourages neighbouring features within every latent vector to store redundant cues about one another, so that the representation remains intelligible even when a subset of features is excised, thereby making the manifold more robust to small latent perturbations..
5 Latent-space properties that facilitate diffusion training
This section embarks on an exploration of the text-latent manifold, asking which of its intrinsic features govern the performance of a diffusion model. All experiments reported here employ autoencoders that compress 128-token texts into 16 latent vectors. The autoencoders are trained on the Wikipedia dataset, whereas the diffusion models are optimised on the smaller yet high-quality ROCStories (mostafazadeh2016corpus, ) dataset. Across all experiments, we generate 1 000 samples and report each metric as the average over these samples. Our experiments spotlight two indispensable qualities — manifold smoothness and resilience to perturbations. The remainder of the section introduces an analysis that quantifies both attributes.
5.1 Smoothness of the latent manifold
Once the autoencoder has been trained, each text $\mathbf{w}$ corresponds to a set of latent vectors $\mathbf{z}$ that can be decoded back into the original text, $\mathbf{w}=D(\mathbf{z})$. Nonetheless, the latent space is far from fully charted: broad regions contain vectors to which no text has ever been assigned, leaving the behaviour of the data density there unknown. Because our diffusion model captures the distribution of texts via the distribution of these latents, yet observes only a finite subset during training, its ability to generalize depends critically on how smoothly that density varies across the manifold. Put differently, the degree to which a locally estimated score function extends to unexplored territory is dictated by manifold smoothness. Inspired by PLANNER (zhang2023planner, ), we investigate this property by conducting the following experiment, which simultaneously evaluates performance at points the model does not encounter during training.
  1. 1.
    Select two random texts from the training corpus and encode them as latent vectors $\mathbf{z}^{(1)},\mathbf{z}^{(2)}~{}\in~{}\mathbb{R}^{N\times d}$.
  2. 2.
    Form a linear interpolation of the endpoints $\mathbf{z}^{\mu}=\mu\,\mathbf{z}^{(1)}+(1-\mu)\,\mathbf{z}^{(2)}$, where $\mu\in[0,1]$ and mid–range $\mu$ values steer $\mathbf{z}^{\mu}$ into regions unseen during training.
  3. 3.
    Apply diffusion noise (Equation 1) to get $\mathbf{z}^{\mu}_{t}$ at different time steps $t$ to measure how unseen regions influence diffusion at different noise levels.
  4. 4.
    Predict a clean latent $\hat{\mathbf{z}}_{0}$ from $\mathbf{z}^{\mu}_{t}$ and decode into text $\mathbf{\hat{w}}$.
  5. 5.
    Evaluate text plausibility with GPT–2 perplexity (PPL), averaged over 1 00010001\,0001 000 random endpoint pairs.
Figure 3: PPL of texts decoded from an interpolation of two latents for Cosmos and CE baseline.
By analysing how perplexity (PPL) evolves with the interpolation coefficient $\mu$ and the diffusion timestep $t$, we obtain a fine-grained picture of manifold smoothness and model robustness far beyond the training support. Figure 3 shows that autoencoder trained only with cross-entropy (CE) loss and our robustness-oriented alternative behave identically when $\mu\!\to\!0$ or $\mu\!\to\!1$, confirming that both models successfully reconstruct train-time input latents. As we move away from the endpoints, perplexity rises for both encoders, but it escalates far more rapidly for the CE baseline — clear evidence that the latent manifold learned by our autoencoder is markedly smoother.
5.2 Reducing the train–inference mismatch
We empirically observe that when a Gaussian diffusion model is trained in the text latent space, a persistent train–inference mismatch often emerges: the latent vector $\hat{\mathbf{z}}$ produced at sampling time can differ markedly from the latent vector the encoder would assign to its own decoded text, $E\!\bigl{(}D(\hat{\mathbf{z}})\bigr{)}\neq\hat{\mathbf{z}}$. This gap has two undesirable effects. First, the decoder becomes unreliable because it must interpret latent vectors contaminated by the diffusion model. Second, the diffusion model repeatedly feeds itself inputs it never saw during training, compounding errors over time. To enable high-quality text generation under this mismatch, both the decoder and the diffusion model must be robust to perturbations in the latent space. In the following sections, we demonstrate through experiments how our training strategy (§4.2.2) fosters such robustness, effectively decreasing train–inference mismatch.
Figure 4: Decoder robustness to latent noising with sequential addition of training modifications.
Decoder robustness.
To assess how well the decoder tolerates perturbations, we inject Gaussian noise into latent vectors of real texts: $\mathbf{z}_{\text{noised}}=\mathbf{z}+\sigma\,\boldsymbol{\varepsilon}$, where $\boldsymbol{\varepsilon}\sim\mathcal{N}(0,I)$, decode the perturbed vector to obtain $\mathbf{\hat{w}}$, and measure its BLEU score against the original text $\mathbf{w}$. Figure 4 shows that two components contribute most to decoder stability: (i) an explicit MSE loss on the decompressor’s output (§4.2.2), which prevents the final-layer norm from exploding, a common effect observed when training solely with cross-entropy loss (gao2021simcse, ), and (ii) latent masking (§4.2.2), which forces the compressor to distribute information more evenly among latent features and remain resilient to partial dropout. Notably, the decoder supplemented with described train-time modifications can reconstruct text almost perfectly even under substantial noise ($\sigma=1$).
Figure 5: Evaluating diffusion model robustness under mid-trajectory noise injection.
Diffusion robustness during generation.
Next, we test the diffusion model’s sensitivity to small perturbations introduced mid-trajectory. For a late generation step $t$, chosen from the second half of the reverse process, where latent representations are already semantically meaningful, we add Gaussian noise with magnitude $\nu$ and continue sampling. The deviation from the original trajectory is quantified by the MSE between the final estimates, $\lVert\hat{\mathbf{z}}-\hat{\mathbf{z}}^{\text{shifted}}\rVert_{2}^{2}$.
As illustrated in Figure 5, the diffusion model trained in our latent space is visibly more stable.
Table 1: Train–inference mismatch for different autoencoder objectives.
Configuration MSE$\bigl{(}\hat{\mathbf{z}},E(D(\hat{\mathbf{z}}))\bigr{)}$
$\mathrm{CE}(w,\hat{w})$ 0.721
[…]
   + Activation aug. 0.582
   + Latent aug. 0.499
Direct mismatch measurement.
Table 1 tracks the train–inference mismatch, quantified as the MSE between $\hat{\mathbf{z}}$ and $E\bigl{(}D(\hat{\mathbf{z}})\bigr{)}$. Each proposed refinement tightens the gap. The consistent drop in error confirms that the enhanced training procedure makes the autoencoder more resilient to diffusion perturbations. Together, these results demonstrate that our training recipe simultaneously strengthens the decoder and the diffusion model, thereby shrinking the train–inference gap.
6 Empirical study of latent representations for diffusion modeling
This section investigates how the design of autoencoder training strategies and the degree of latent space compression influence the quality of text generation via diffusion.
6.1 Empirical analysis of autoencoder training regimes
In this part of our study we assess the contribution of each proposed latent space augmentation to the final quality of the diffusion model. For the ablation study, we compress text from 128 BERT representations down to 16 latent vectors. All autoencoders are trained on the Wikipedia dataset, while the diffusion models are trained on the ROCStories (mostafazadeh2016corpus, ) dataset.
Table 2: Comparison of autoencoder training regimes on text generation quality. Features are added cumulatively from top to bottom. The gray-highlighted row indicates the selected configuration used in the final model.
Configuration MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$
CE($\bf{w,\hat{w}}$) $0.294_{.036}$ $71.5_{.9}$ $0.298_{.001}$
+ MSE($\bf{h,\hat{h}}$) $0.464_{.024}$ $57.1_{.7}$ $0.344_{.002}$
+ Random masking (rate)
   0.1 $0.465_{.029}$ $54.5_{.4}$ $0.356_{.003}$
   0.2 $0.565_{.023}$ $50.5_{.5}$ $0.342_{.001}$
   0.3 $0.586_{.015}$ $42.8_{.7}$ $0.335_{.003}$
   0.4 $0.514_{.011}$ $46.4_{.3}$ $0.344_{.002}$
   0.5 $0.461_{.025}$ $51.8_{.7}$ $0.362_{.005}$
+ Gaussian noising ($\delta$)
   0.5 $0.679_{.01}$ $36.8_{.3}$ $0.310_{.003}$
   0.6 $0.724_{.015}$ $35.6_{.7}$ $0.324_{.004}$
   0.7 $0.767_{.011}$ $33.6_{.3}$ $0.328_{.002}$
   0.8 $0.725_{.018}$ $36.4_{.7}$ $0.333_{.002}$
   0.9 $0.664_{.023}$ $44.3_{.8}$ $0.294_{.003}$
+ Latent dropout (rate)
   0.1 $0.771_{.014}$ $33.9_{.7}$ $0.325_{.004}$
   0.2 $0.781_{.029}$ $32.6_{.8}$ $0.322_{.004}$
   0.3 $0.812_{.012}$ $31.9_{.5}$ $0.328_{.003}$
   0.4 $0.836_{.009}$ $30.2_{.5}$ $0.322_{.004}$
   0.5 $0.724_{.012}$ $34.8_{.4}$ $0.320_{.002}$
Evaluation metrics.
We evaluate the unconditional generation capacity of models with three complementary metrics that, together, capture textual quality, lexical diversity, and distributional fidelity. First, we report perplexity (PPL), computed with GPT-2 Large (radford2019language, ). Second, we quantify diversity using the score $\mathrm{\textbf{div}}(y)=\prod_{n=2}^{4}\tfrac{\#\,\text{unique }n\text{-gramsin }y}{\#\,n\text{-grams in }y}$, where $y$ denotes the set of generated texts. Finally, because low perplexity can be achieved by simple or repetitive texts, we also compute the MAUVE score (pillutla2021mauve, ), which measures the distributional alignment between generated and reference texts and thus offers a broader view of generation quality. For every model we generate $1{,}000$ samples and, when calculating MAUVE, draw an equally sized reference subset from the held-out test split. The entire evaluation procedure is repeated ten times with different random seeds, and we report the mean and standard deviation of the resulting scores across runs.
Results.
Table 2 demonstrates that each proposed feature contributes significantly to the model performance. Remarkably, adding MSE penalty between the predicted and reference BERT activations alone boosts all quality metrics by about 50%. Subsequent augmentation of the BERT activations yields a further leap, underscoring the value of intermediate contextual, semantically grounded representations. This stage rises MAUVE to 0.767, already achieving the performance obtained when the diffusion model is trained on the full-length, uncompressed BERT representations (0.767 vs. 0.762 for TEncDM (shabalin2025tencdm, )). Finally, dropping random features from latent vectors pushes MAUVE beyond 0.83, lowers perplexity to 30.2, keeping diversity essentially unchanged.
These results provide a clear answer to the question posed in the introduction of this study: text representations can indeed be mapped into a more compact latent space, where a trained diffusion model performs comparably to traditional token-level counterparts without sacrificing the text generation quality. For all subsequent autoencoders we lock in the configuration highlighted in Table 2: the MSE penalty is kept, the random-masking rate is fixed at $0.3$, Gaussian noising is applied with $\delta=0.7$, and latent masking rate is set to $0.4$.
Additionally, we validate all proposed autoencoder configurations by training the diffusion model on the larger Wikipedia dataset. Detailed results of this evaluation are presented in Section C.2. Furthermore, we explore the use of a variational prior in the latent space and observe that it offers no clear advantage. A thorough analysis is provided in Section C.1.
6.2 Impact of scaling the number of latent vectors
Figure 6: Impact of scaling $N$ on diffusion generation quality.
In this section, we examine how diffusion generation quality varies when varying the number of latent vectors $N$. We keep the training pipeline fixed according to the hyperparameters detailed in Section 6.1. We do not modify the embedding dimension $d$, because it would necessitate architectural modifications to the diffusion model, making it difficult to isolate the impact of latent space diffusability from changes in model capacity and structure.
Table 3: Impact of scaling the number of latent vectors $N$ on unconditional generation quality.
$N$ MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$
Source 0.953 21.7 0.403
BERT repr. 0.762.043 29.1.9 0.295.002
2 0.172.018 109.5.9 0.344.004
4 0.197.016 70.1.8 0.354.003
8 0.403.007 51.9.7 0.327.002
16 0.836.009 30.2.5 0.322.004
32 0.901.008 27.3.4 0.346.004
64 0.923.009 26.7.2 0.347.003
128 0.940.011 26.3.4 0.346.004
Figure 6 compares baseline cross-entropy compressor to the robustness-oriented compressor introduced in Section 4.2.2. We observe that for CE compressor decrease of $N$ leads to a rapid degradation in quality. In contrast, the proposed autoencoder maintains high generation quality even under substantial compression. It achieves an 8×8$\times$8 × reduction in latent sequence length while surpassing the quality of uncompressed representations.
Table 8 provides further insights by comparing generation quality across different numbers of latent vectors $N$. Although configuration $N=16$ matches the quality of BERT representations baseline, we observe a notable quality gap compared to the configuration with $N{=}128$: the MAUVE score decreases from $0.940$ to $0.836$, while perplexity increases from $26.3$ to $30.2$. However, adopting a less aggressive compression to $N{=}32$ (a 4×4$\times$4 × compression) results in only minor quality degradation, still significantly outperforming the baseline. Based on these observations, together with results from Section 7, we recommend limiting latent compression to approximately 4×4$\times$4 × to optimally balance quality and computational efficiency.
7 Comparison across generative paradigms
We evaluate the performance of Cosmos on four distinct text generation tasks: unconditional story generation (ROCStories (mostafazadeh2016corpus, )), text summarization (XSum (narayan2018don, )), detoxification (ParaDetox (logacheva-etal-2022-paradetox, )), and question generation (SQuAD2.0 (minaee2021deep, )). For unconditional story generation we employ metrics discussed in Section 6.1. For text summarization and question generation, we report BERTScore (BS) (zhang2019bertscore, ). The detoxification task is evaluated using BLEU-4 (BLEU). We further assess models’ efficiency by comparing inference time for unconditional generation on sequences with 128 and 512 tokens. Additional metrics and implementation details are provided in the Section C.3.
Baselines.
We benchmark our proposed method against a diverse set of generative paradigms. These include Gaussian diffusion on embeddings (DiffuSeq (gong2022diffuseq, ), SeqDiffuSeq (yuan2022seqdiffuseq, ), AR-Diffusion (wu2023ar, )), simplex-based diffusion (TESS (mahabadi2023tess, )), masked diffusion (SEDD (lou2023discrete, )), alternative latent diffusion baselines (LD4LG (lovelace2023latent, ), TEncDM (shabalin2025tencdm, )), and autoregressive models (GPT-2 (radford2019language, ) and GPT-Neo (gpt-neo, )). All models operate within a comparable parameter scale, ranging from approximately 100M to 200M. To ensure fairness, all baselines were faithfully reimplemented and trained on generative tasks using training protocols closely aligned with those described by their original authors.
Autoencoder setup across tasks.
[…]
Table 4: Comparison with autoregressive and diffusion baselines across four generative tasks. Inference time (in seconds) is reported for sequence lengths of 128 and 512 tokens. The best-performing scores are shown in bold, while the second-best scores are underlined.
ROCStories XSum ParaDetox SQuAD2.0 Time (s)
Method MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$ BS $\uparrow$ BLEU $\uparrow$ BS $\uparrow$ $L=128$ $L=512$
Source text 0.953 21.7 0.403
GPT2 0.789 20.5 0.252 0.690 0.677 0.680 $\mathbf{1.2_{.1}}$ $42.8_{.1}$
GPT Neo 0.720 19.9 0.258 0.621 0.610 0.665 $2.6_{.1}$ $45.2_{.2}$
AR-Diffusion 0.066 41.8 0.101 0.568 0.647 0.569 $226.4_{.2}$ $602_{2.1}$
DiffuSeq 0.086 50.5 0.124 0.588 0.679 0.563 $215.9_{1.2}$ $1565_{.7}$
SeqDiffuSeq 0.103 29.3 0.137 0.617 0.688 0.574 $100_{.5}$ $601_{.5}$
TESS 0.061 22.4 0.163 0.627 0.693 0.667 $1441_{.2}$ $5984_{.0}$
SEDD 0.598 70.8 0.336 0.576 0.666 0.443 $15.0_{1.1}$ $60.3_{1.0}$
LD4LG 0.716 30.6 0.331 0.702 0.708 0.641 $27.9_{.5}$ $102_{.2}$
TEncDM 0.762 29.1 0.295 0.699 0.619 0.703 $29.6_{.1}$ $180_{.3}$
CosmosN=16 0.836 30.2 0.322 0.654 $5.8_{.1}$
CosmosN=128 0.940 26.3 0.346 0.704 0.694 0.708 $35.1_{.1}$ $\mathbf{36.6_{.1}}$
Results.
[…]
8 Conclusion
In this paper, we present Cosmos, a novel approach to text generation that shifts the focus from high-dimensional token-level representations to a compact latent space tailored for diffusion modeling. By learning a compressed and smooth latent representation through carefully designed autoencoder objectives, Cosmos enables high-quality generation while reducing inference time. Our empirical results show that, Cosmos matches or outperforms traditional token-level diffusion and autoregressive baselines across diverse tasks. Our findings challenge the dominance of token-level models and highlight latent diffusion as a promising direction for building fast, high-quality language models.
Appendix
[…]
section0em2em
Appendix A Dataset descriptions
ROCStories
This dataset [23] consists of five-sentence stories and serves as a well-established, small-scale, unconditional benchmark in language diffusion research. It contains a total of 98,161 instances, of which 88,161 are used for training, 10,000 for validation.
Wikipedia
For large-scale experiments, we use the English Wikipedia subset from the ROOTS corpus [14]. The resulting dataset comprises 8.6 million sequences, with 50,000 held out for validation.
XSum
This dataset [25] is used for abstractive summarization and comprises 204,000 BBC news articles paired with summaries covering diverse topics (e.g., sports, politics). It includes 204,045 training instances, 11,332 validation instances, and 11,334 test instances. We truncate input articles to 512 tokens and limit reference summaries to 50 tokens.
SQuAD2.0
This question-answering dataset [30] requires identifying the text span that answers a given question or indicating that no answer is possible from the provided context. For generative modeling, we reverse the task: given a context and an answer, the model generates the corresponding question. The dataset contains 130,319 training and 11,873 test instances. Contexts are truncated to 512 tokens.
ParaDetox
For small-scale conditional generation experiments, we use the ParaDetox dataset [16], which comprises 19,766 pairs of toxic and neutral comments. Both context and target sequences are truncated to 40 tokens.
Appendix B Implementation details
B.1 Latent diffusion pipeline
With a diffusable latent manifold in hand (§4.2.2), we now describe the diffusion model that operates entirely within this space. Our setup follows Eq. (2): we train a denoiser that receives a noisy latent $\mathbf{z}_{t}$ at diffusion time $t\!\in\![0,1]$ and predicts the corresponding clean latent $\mathbf{z}_{0}$. The loss is the simple mean-squared error $\lVert\mathbf{z}_{0}-z_{\theta}(z_{t},t)\rVert_{2}^{2}$, where $z_{\theta}$ shares parameters across all time-steps. In line with earlier work [18, 33], we employ self-conditioning: on $50\%$ of training updates the model is fed its own previous estimate of $\mathbf{z}_{0}$ as an additional input. This iterative refinement enables the model, at inference time, to reuse its past predictions, producing markedly crisper and more coherent text. We adopt the noise schedule introduced by TEncDM [33]: $\alpha_{t}=\frac{1}{1+\tan(t\pi/2)^{2}\cdot d^{2}}$, where parameter $d$ controls he rate at which noise is injected during the diffusion steps. At inference time we use Euler solver with 200 uniform steps. The denoiser itself inherits the TEncDM [33] architecture. It is a 12-layer Transformer with 12 attention heads and a hidden size of 768. In total, the network comprises roughly $\approx\!130$M parameters.
Conditional generation.
For sequence-to-sequence generation, we train a diffusion model in a conditional setting. The model learns to reconstruct a noisy latent representation of the target text, conditioned on the latent representation of the source text generated by our autoencoder. Consequently, the source text’s vector representation is also leveraged in a compressed format, enhancing both training and inference efficiency. The diffusion model is conditioned on the source latent, which is first processed by a 12-layer transformer encoder. This encoder extracts relevant features and derives a suitable conditioning representation from the source latent. These encoded source representations are then injected into the diffusion model using cross-attention mechanisms. For conditional tasks, we initiate the training of our diffusion model by initializing its denoiser weights with those pretrained on an unconditional task.
B.2 Training and inference configuration
[…]
ROCStories Wikipedia XSum SQuAD2.0 Paradetox
Optimizer AdamW
Learning Rate 2e-4
($\beta_{1}$, $\beta_{2}$) (0.9, 0.98)
Warmup Steps 1000
Learning Rate Schedule Constant
Weight Decay 0.01
Gradient Clipping 1
EMA Decay 0.9999
Batch Size 1024
Training Steps 200k 500k 100k 100k 10k
Max Seq Length 80 512 64 64 40
Max Context Length 512 512 40
Sampling steps 200
Schedule parameter 5 3 7 7 9
Autoencoder Training Time 14h
Diffusion Training Time 1d 5h 2d 12h 1d 1h 1d 9h 14h
Table 5: Training details for Cosmos across different datasets.
B.3 Inference time benchmarking configuration
We measure the generation time for a batch of 512 samples per model on a single NVIDIA A100 GPU using the bfloat16 data type. Inference settings for all baseline models follow the default configurations of their respective repositories, except for SEDD, where we adopt 32 diffusion steps in accordance with the original paper’s setup for matching autoregressive quality. We report the mean and standard deviation across five independent runs.
Appendix C Additional autoencoder training analysis
C.1 Marginal utility of variational prior in text autoencoder
In this section, we explain why incorporating a variational prior does not significantly enhance the performance of our text autoencoder.
In a classical Variational Autoencoder (VAE), the encoder maps each input to the parameters of a Gaussian latent distribution, i.e., a mean vector $\mu$ and a (diagonal) variance vector $\sigma^{2}$. Training minimises a reconstruction term together with the Kullback–Leibler (KL) divergence between the encoder distribution and an isotropic prior $\mathcal{N}(0,I)$. For a latent matrix $z\in\mathbb{R}^{N\times d}$, the KL term is
$D_{\mathrm{KL}}\!\bigl{(}\mathcal{N}(z;\mu,\sigma^{2})\,\|\,\mathcal{N}(0,I)\bigr{)}=\tfrac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{d}\!\bigl{(}\mu_{ij}^{2}+\sigma_{ij}^{2}-\log\sigma_{ij}^{2}-1\bigr{)}.$
Our text autoencoder is trained with the following objective:
$\mathcal{L}=\mathrm{CE}(w,\hat{w})\;+\;\mathrm{MSE}(h,\hat{h})\;+\;\beta\,D_{\mathrm{KL}}\!\bigl{(}\mathcal{N}(z;\mu,\sigma^{2})\,\|\,\mathcal{N}(0,I)\bigr{)},$
where $\mathrm{CE}(w,\hat{w})$ is the token-level cross-entropy, $\mathrm{MSE}(h,\hat{h})$ matches reconstructed contextual representations, and $\beta\!\geq\!0$ balances reconstruction fidelity against latent regularisation.
During training, the decoder receives a stochastic latent sample
$z_{s}=\mu+\sigma\odot\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,I),$
so that the KL term nudges $(\mu,\sigma)$ towards the prior. Increasing $\beta$ strengthens this pressure, producing a smoother latent geometry at the cost of higher reconstruction error; decreasing $\beta$ does the opposite as encoder tends to push latents away from each other.
In our experiments, we observed that the performance is highly sensitive to the choice of  $\beta$. Table 6 presents the results for different values of  $\beta$.
Table 6: Impact of the KL weight $\beta$ on diffusion generation quality. Arrows indicate the preferred direction.
$\beta$ MAUVE $\uparrow$ PPL $\downarrow$ DIV $\uparrow$
0 0.767 33.6 0.328
0.0001 0.733 39.6 0.329
0.001 0.764 33.3 0.328
0.01 0.765 32.6 0.326
0.1 0.658 39.6 0.329
1 0.011 677.1 0.587
The best perplexity is achieved at $\beta=0.01$, but the gain over the baseline ($\beta=0$) is marginal (33.6 $\rightarrow$ 32.6). For larger $\beta$ the model collapses: at $\beta=1$, perplexity explodes and MAUVE falls.
This limited improvement can be attributed to the decoder’s robustness to Gaussian noise in the latent space. As illustrated in Figure 4, the decoder maintains reasonable performance even when sampling from the prior with $\sigma=1$, indicating that the latent space is sufficiently robust without strong KL regularization.
Alternative regularisers, notably latent masking, yield larger and more robust improvements. We therefore omit the KL term in the final model and rely solely on latent masking to shape the representation space.
C.2 Empirical analysis of autoencoder training regimes on Wikipedia dataset
Table 7: Comparison of autoencoder training regimes on text generation quality on Wikipedia. Features are added cumulatively from top to bottom.
Configuration MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$
CE($\bf{w,\hat{w}}$) $0.056$ $385.6$ $0.590$
+ MSE($\bf{h,\hat{h}}$) $0.069$ $344.4$ $0.625$
+ Random masking (rate)
   0.3 $0.081$ $289.1$ $0.594$
+ Gaussian noising ($\delta$)
   0.7 $0.103$ $205.0$ $0.588$
+ Latent dropout (rate)
   0.4 $0.112$ $186.7$ $0.581$
We also conduct a large-scale ablation on the Wikipedia dataset with $N=16$ latents and max sequence length $L=128$, following a setup analogous to that in section 6.1. Results, presented in Table 7, reveal trends consistent with those observed in the small-scale ablation. Notably, applying the proposed training modifications leads to substantial improvements in evaluation metrics. The final model approaches the performance of the TEncDM baseline [33], while operating in a latent space that is 8×8$\times$8 × smaller.
Additionally, we investigate how scaling the latent space length on Wikipedia affects performance. We experiment with latent sequence lengths ranging from $16$ to $128$. As expected, we observe steady performance improvements with longer latent sequences. Results are provided in Table 8.
Table 8: Impact of scaling the number of latent vectors $N$ on unconditional generation quality on Wikipedia.
$N$ MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$
Source 0.953 21.7 0.403
BERT repr. 0.109 173.1 0.562
16 0.112 186.7 0.581
32 0.149 134.4 0.579
64 0.157 128.1 0.582
128 0.173 118.2 0.592
C.3 Additional details of comparison across generative paradigms
Table 9: Comparison with autoregressive and diffusion baselines across three generative tasks. The best-performing scores are shown in bold, while the second-best scores are underlined.
ParaDetox Xsum SQuAD2.0
Method BLEU $\uparrow$ J-Score  $\downarrow$ R-1/2/L  $\uparrow$ BS  $\uparrow$ R-L  $\uparrow$ BS  $\uparrow$
GPT2 0.677 0.604 0.283/0.082/0.218 0.690 0.332 0.680
GPT Neo 0.610 0.492 0.231/0.045/0.171 0.621 0.245 0.665
AR-Diffusion 0.647 0.465 0.268/0.059/0.206 0.568 0.185 0.569
DiffuSeq 0.679 0.475 0.189/0.130/0.136 0.588 0.186 0.563
SeqDiffuSeq 0.688 0.486 0.286/0.067/0.213 0.617 0.194 0.574
TESS 0.694 0.587 0.317/0.116/0.264 0.627 0.339 0.667
SEDD 0.666 0.001 0.200/0.033/0.138 0.576 0.086 0.443
LD4LG 0.708 0.580 0.303/0.100/0.246 0.702 0.211 0.641
TEncDM 0.619 0.496 0.319/0.107/0.253 0.699 0.323 0.703
CosmosN=16 0.649 0.497
CosmosN=128 0.694 0.554 0.328/0.114/0.258 0.704 0.339 0.708
We also present extended results for our conditional generation tasks using a broader variety of evaluation metrics. For XSum and SQuAD2.0, we report ROUGE scores [15], a standard metric that evaluates text quality based on n-gram overlap with reference outputs. For ParaDetox, we additionally include the J-score, which is defined as the product of style accuracy, fluency, and content preservation. Extended results are summarized in Table 9.
Appendix D Limitations
There are several limitations to our study that point to promising directions for future work. First, jointly training the autoencoder and diffusion model remains an open research direction. Such an approach may significantly improve training efficiency. Second, the latent dimensionality $d$ is held fixed in all experiments, because changing $d$ would necessitate redesigning diffusion backbone; a systematic sweep of this hyper-parameter is therefore left to dedicated follow-up work. Finally, we report results with relatively small backbones of roughly 130M parameters to keep ablations rapid and compute budgets fair; scaling the architecture is an orthogonal engineering effort that is expected to reinforce the empirical trends observed here.
Appendix E Societal Impact
Our research introduces an alternative modeling approach for language, centered on latent continuous diffusion. We hold the view that this method does not bring about significant new societal risks exceeding those already connected with current language models.
Appendix F Generation examples
To give a qualitative sense of model behaviour, we first present representative unconditional generations: Table 10 shows random samples from CosmosN=128 and the TEncDM baseline on ROCStories, while Table 11 does the same for Wikipedia. We then move to the conditional setting, providing sequence-to-sequence outputs for three benchmarks: summaries for XSum in Table 12, detoxified rewrites for ParaDetox in Table 14, and question-generation examples for SQuAD 2.0 in Table 13.
Table 10: Randomly generated samples for ROCStories dataset.
TEncDM CosmosN=128
Amy wanted to buy a new dress for the dance! She shopped around for ten minutes. She tried on a dress that was too big. Amy loved that the dress was too grab for her size. She found that it was a little too long to wear! Amy and her friend Sue were excited about sixth grade together. The three of them visited Amy’s online shoe store. Amy’s friend Beth arrived at the shoe store and Sue tried some dresses. The girls checked out and looked at all the collection. Amy felt like she was being an idiot.
Maggie was a nasty girl in school. When she moved to a small town, she had no Mae friends. After school, sheared out and made no friends. Sometimes they called her back, and they left her alone. She finally understood why she needed to be friends. Jay was sitting at home bored. He was looking for an activity to keep him active. He decided to play a game of basketball. He played basketball for an hour. He was able to have a fun time that afternoon.
My car was becoming very worn out. I went to my mechanic to get checked out. He told me that I had a flat tire. I went to NCaven to get it fixed. I still made it to work in no time. Sam was working late. He didn’t have money to pay all his bills. He was having trouble getting paid. Sam decided to quit his work job. He was able to get another job to pay his bills.
My daughter is moving to a new school next year. She was nervous about moving to this new location. I am afraid of a leaker of all of her relationships. We figured it out and made some alterations on the team. This is where she will be graduating high school. Debra hasn’t written a check in weeks. She’s been very stressed and frustrated with paying bills. Last week, she acted horribly during her check. The bank quickly ran a check - up. To her shock, she found out that she owed nearly a million in cash.
John went to the library. The librarian told him to read more books. John went to the booktore. He tried to read 6 books. He couldn’t decide which to buy. Tina never thought she needed her son to be cool enough to swim. So she begged him to do so while he was supposed to. So he practiced and swam all the time. But when it was time to go outside, he constantly got sore. Tina was too stressed to let him swim for 3 hours.
Table 11: Randomly generated samples for the Wikipedia dataset using CosmosN=128.
The press said working conditions for the match did not have to be finalised. Even the organizers had planned for the date of the match to be 19 August 2017 to accommodate a crowd of some 500, 000 people watching the match.
Bob Hall served as Foundation founding president from 1990 to 1991. In 1993, Hansen was elected President of the College of Bymphoschipelgeons. Upon his retirement from this position, he also served as the chairman of the Scientific Advisory Committee of the College of Medical Surgeonsgeons ( IMS ). In 1996, publisher magazine selected Bob Hall as his candidate for the next U. S. presidential nomination for 2000 and 2001. He died October 26, 2013, at his home in Tampa, Florida from the causes of kidney failure at the age of 71, caused by a cerebral kidney failure in 1980. He is buried at the
He was excluded from the committee because Johnson refused to approve the text of the bill, and two senators refused to attend discussing the legal provisionss of the Seventh Amendment, but invited senior officials to hear it. Both the Congress House and the Weed Hermans challenged the committee’s request to have the Seventh Amendment be argued in the Supreme Court’s decision during the subsequent legislative hearings. Both attributed the premise premisel of Johnson’s Amendment to internal law, which allocated the Congress to adopt executive decrees beginning in January 1699. The law, for its part, was however applied by the majority of the
Folk music typically contains a mix of German, Russian, Russian, Iranian, Egyptian, Norwegian, Turkish, Jewish and Tabad music, though typically the songs are from a specific religion or background. Icelandmar0̆0edk, for instance, distinguish genderly distinguishes folk music from the multi - diverse genres, primarily rock and jazz.. Folk music is consequently not subegoaticly distinct from rock, which in turn, which consists primarily of music of folk including jazz, blues, nor rock, and industrial rock. Due to the diversity of genres in these genres, certain new genres of music also exist. Pandit music, such as
In 1974, Parker became the Director of the Office of Virginia’s Chamber of Trade and Commerce, a position he would hold under increased pressure from his fellow Democratic opponent. He oversaw the J. Howard administration, which lasted for 11 years headed by William J. P honorter.
Table 12: Samples from XSum summarization dataset. Parts of the articles are omitted for brevity.
Article: Two-year-old Lane Thomas Graves had been playing in the sand near the resort’s Seven Seas Lagoon when he was dragged underwater by the creature… The lighthouse has been installed near to where the attack occurred… A Disney spokesperson said they hoped the monument would spread awareness for the Lane Thomas Foundation, which also uses the lighthouse as its logo. Who is liable for alligator boy’s death? "The lighthouse sculpture has been installed to help spread awareness of the Lane Thomas Foundation, which was established to provide assistance and support to families whose children need organ transplants," Walt Disney World said in a statement.
Reference: Walt Disney World has unveiled a lighthouse memorial for a young boy who was killed by an alligator while on holiday at the Florida theme park.
TEncDM CosmosN=128
A statue of a Florida boy who died after being rescued from a beachquarina during water has been honoured by US television provider Disney. Walt Disney Disney has unveiled a lighthouse statue in memory of a young boy who died when he was stabbed by an alligator at a Florida resort in Florida.
Article: The Sky Blues currently play in Coventry’s Ricoh Arena but had a long dispute with the stadium’s previous owners… In a statement, Rugby Borough Council said its leader and the council’s executive director and head of planning had met with Coventry City in March. "The club requested the meeting to understand how the council would deal with any planning application for potential stadium sites in the borough of Rugby," it said. It said the plans would need to be finalised by September to be included in the council’s local plan, but added that a site had yet to be identified. Peter Ward, from Sky Blues Supporters’ Consultative Group, said he was pleased to hear that things were "moving" with the club’s search for a new home. "It’s good that finally there is some evidence things
Reference: Planners in Rugby have revealed they have been in talks with Coventry City Football Club about building a stadium in the borough.
TEncDM CosmosN=128
Coventry City Council says it is looking on whether potential plans for a 0̆0a33m Super League stadium in Coventry. Coventry City fans say they will meet with the city council over a proposed move from the club club to a new stadium.
Table 13: Samples from SQuAD2.0 question generation dataset.
Context: Temporal measurement has occupied scientists and technologists, and was a prime motivation in navigation and astronomy. Periodic events and periodic motion have long served as standards for units of time. Examples include the apparent motion of the sun across the sky, the phases of the moon, the swing of a pendulum, and the beat of a heart. Currently, the international unit of time, the second, is defined by measuring the electronic transition frequency of caesium atoms (see below). Time is also of significant social importance, having economic value (ẗime is money)̈ as well as personal value, due to an awareness of the limited time in each day and in human life spans.
Answer: Temporal measurement.
Question: What has been a prime motivation in astronomy and navigation?
TEncDM CosmosN=128
What has a significant role in surveying units of time? What have technologists filled time in astronomy?
Context: In Book 11 of his Confessions, St. Augustine of Hippo ruminates on the nature of time, asking, Ẅhat then is time? If no one asks me, I know: if I wish to explain it to one that asketh, I know not.Ḧe begins to define time by what it is not rather than what it is, an approach similar to that taken in other negative definitions. However, Augustine ends up calling time a 2̆01cdistention2̆01d of the mind (Confessions 11.26) by which we simultaneously grasp the past in memory, the present by attention, and the future by expectation.
Answer: St. Augustine of Hippo.
Question: Who commented on the nature of time in Book 11 of his confessions?.
TEncDM CosmosN=128
Who wrote " " nature of timefessions ? Which philosopher ruminates about explaining the nature of time?
Table 14: Random samples for the detoxification task on the ParaDetox dataset.
Input: fucking imagine obama just put the hands up and keep a good distance between he and cruz .
Reference: Imagine Obama just put the hands up and distanced himself with Cruz.
TEncDM CosmosN=128
2̆018 obama just put the hands up and keep a good distance between him and cruz. imagine obama just put the hands up and keep a good distance between he and cruz.
Input: you just summed up how fucking stupid politics is in one comment
Reference: You just summed up how politics is in one comment.
TEncDM CosmosN=128
you just sum up up how bad politics is in one comment you just sum up up how how stupid politics is in one comment
Input: only churches give a crap it seems.
Reference: only churches care it seems.
TEncDM CosmosN=128
Only churches care bad it it seems.. only churches give a thing it seems.
Input: second , maybe if you douches would read the damn edit summaries , this wouldnt happen .
Reference: second , maybe if you would read the edit summaries , this wouldnt happen .
TEncDM CosmosN=128
2̆01d, maybe if you would read the editmaries. If you read the edit summaries, this wouldn’t happen.
Input: what we witnessed here today was a fucking miracle, and i want you to acknowledge that.
Reference: What we witnessed here today was a miracle, and I want you to acknowledge that.
TEncDM CosmosN=128
What we witnessed here today was a miracle, and I want you to acknowledge that. what we witnessed here today was a a miracle, and i want to acknowledge that.

### v2

@@
 We introduce Cosmos, a novel approach to text generation that operates entirely in a compressed, smooth latent space tailored specifically for diffusion.
-This space is learned using an autoencoder trained simultaneously for token-level reconstruction and alignment with frozen activations from a pretrained language encoder, providing robust semantic grounding and enabling effective perturbation-based augmentations.
-Empirically, we demonstrate that text representations can be compressed by 8×8$\times$8 × while maintaining generation quality comparable to token-level diffusion models.
-Furthermore, increasing the latent sequence length allows Cosmos to surpass both diffusion-based and autoregressive baselines.
+This space is learned using an autoencoder trained simultaneously for token-level reconstruction and alignment with frozen activations from a pretrained language encoder, providing robust semantic grounding and enabling effective perturbation‑based augmentations.
+Empirically, we demonstrate that text representations can be compressed up to $8\times$ while maintaining generation quality comparable to token‑level diffusion models.
+Furthermore, increasing the latent sequence length allows Cosmos to surpass both diffusion‑based and autoregressive baselines.
 We evaluate Cosmos on four diverse generative tasks including story generation, question generation, summarization, and detoxification and compare it with various generative paradigms.
-Cosmos achieves comparable or superior generation quality while offering more than 2×2$\times$2 × faster inference.
+Cosmos achieves comparable or superior generation quality while offering more than $2\times$ faster inference.
+Code is released at GitHub.
+11footnotetext: Corresponding author: vmeshchaninov@hse.ru
 1 Introduction
@@
 We hypothesize that the token-level encoding used in contemporary language models is heavily overparameterized for the purpose of sequence generation.
-Building on the successes of latent diffusion in vision, we pose the following question: How far can we compress textual information into a compact latent space while still matching or even exceeding the generative fidelity of conventional token-level representations?
+Building on the successes of latent diffusion in vision, we pose the following question: How far can we compress textual information into a compact latent space while still matching or even exceeding the generative fidelity of conventional token‑level representations?
 To answer this question, we develop an autoencoder that maps text into a lower-dimensional space and then train a diffusion model directly in that space.
@@
 Our key contributions are as follows:
-  • We propose Cosmos — a training recipe for a COmpressed and SMOoth latent Space, that allows to train a diffusion model in more compact latent space while matching the quality of token-level diffusion baselines.
-  • We demonstrate that modestly scaling the number of latent vectors enables latent-space diffusion to surpass both token-level diffusion and autoregressive models on an unconditional text generation task.
-  • On standard text-generation benchmarks, the proposed latent diffusion achieves up to $\mathbf{2\times}$ faster sampling than conventional token-level diffusion models, while matching or slightly exceeding them in quality and diversity.
+  • We propose Cosmos — a training recipe for a COmpressed and SMOoth latent Space, that allows to train a diffusion model in more compact latent space while matching the quality of token‑level diffusion baselines.
+  • We demonstrate that modestly scaling the number of latent vectors enables latent‑space diffusion to surpass both token‑level diffusion and autoregressive models on an unconditional text generation task.
+  • On standard text‑generation benchmarks, the proposed latent diffusion achieves up to $\mathbf{2\times}$ faster sampling than conventional token‑level diffusion models, while matching or slightly exceeding them in quality and diversity.
 2 Related work
@@
 We devise a training procedure that shrinks BERT-level representations by up to $\bf{8\times}$ while preserving, and in some cases enhancing, their suitability for latent-space diffusion.
+Our study builds upon the approaches reviewed here. For additional context and a discussion of related topics please see Appendix B.
 3 Preliminary
@@
 Latent diffusion model.
-We train a Gaussian diffusion model is space formed by $\mathbf{z}$, following Eqs. (1)–(2). Because $N$ is small, the diffusion model runs faster in this space.
+We train a Gaussian diffusion model is space formed by $\mathbf{z}$, following Eqs. (1)–(2). Because NN is small, the diffusion model runs faster in this space.
 Decompressor.
@@
 4.2 Learning a compact text latent space
-The previous Section 4.1 outlined our end-to-end pipeline: a frozen contextual encoder supplies token representations, a compressor distils them into a fixed-length latent matrix, and a diffusion model operates solely in this latent space.
+The previous Section˜4.1 outlined our end‑to‑end pipeline: a frozen contextual encoder supplies token representations, a compressor distils them into a fixed‑length latent matrix, and a diffusion model operates solely in this latent space.
 We now zoom in on the compression and decompression stages.
@@
 4.2.1 How compact can text latents be?
-Figure 2: Token-level reconstruction accuracy on Wikipedia (512 tokens) as a function of $N$.
-Figure 2 reports token-level reconstruction accuracy on Wikipedia dataset as a function of the latent sequence length $N$.
-The compressor outlined in §4.1 is trained using a token-level cross-entropy objective.
-Remarkably, a lightweight compressor equipped with only 12 transformer blocks is able to encode 512-token sequences into only $N=16$ latents with $100\%$ reconstruction accuracy, achieving a 32×32$\times$32 × compression relative to initial 512 hidden states produced by BERT.
-These results paint an encouraging picture: if such a compact latent manifold can be modeled generatively, text generation could proceed in a space whose dimensionality is much smaller than that of the original token embeddings, promising dramatic speed-ups.
+Figure 2: Token‑level reconstruction accuracy on Wikipedia (512 tokens) as a function of $N$.
+Figure˜2 reports token‑level reconstruction accuracy on Wikipedia dataset as a function of the latent sequence length $N$.
+The compressor outlined in §4.1 is trained using a token‑level cross‑entropy objective.
+Remarkably, a lightweight compressor equipped with only 12 transformer blocks is able to encode 512‑token sequences into only $N=16$ latents with $100\%$ reconstruction accuracy, achieving a $32\times$ compression relative to initial 512 hidden states produced by BERT.
+These results paint an encouraging picture: if such a compact latent manifold can be modeled generatively, text generation could proceed in a space whose dimensionality is much smaller than that of the original token embeddings, promising dramatic speed‑ups.
 However, as we show further, high reconstruction accuracy by itself does not imply that the latent manifold is suitable for generative modeling.
 4.2.2 Learning robust compact text representations
-The reconstruction study (§4.2.1) reveals that aggressive text compression can be loss-free.
+The reconstruction study (§4.2.1) reveals that aggressive text compression can be loss‑free.
 However, for a robust diffusion generation compressed space should satisfy additional requirements.
-Empirically, we observe that if the latent manifold lacks smoothness and robustness (Section 5), a Gaussian diffusion model fails to sample latents that yield high-quality texts.
+Empirically, we observe that if the latent manifold lacks smoothness and robustness (Section˜5), a Gaussian diffusion model fails to sample latents that yield high‑quality texts.
 To improve diffusability of the latent space, we employ three complementary strategies to the autoencoder training.
-Figure 1 provides an overview of the proposed approach.
+Figure˜1 provides an overview of the proposed approach.
 MSE regularisation on encoder activations.
-Alongside the cross-entropy loss between the original tokens and their reconstructions, we add a mean-squared error penalty between the frozen text encoder outputs $\mathbf{h}$ and their reconstructions $\mathbf{\hat{h}}$.
+Alongside the cross‑entropy loss between the original tokens and their reconstructions, we add a mean‑squared error penalty between the frozen text encoder outputs $\mathbf{h}$ and their reconstructions $\mathbf{\hat{h}}$.
 This auxiliary objective forces the compressor to preserve the semantics carried by the contextual representations $\mathbf{h}$.
-Activation-space perturbations.
+Activation‑space perturbations.
 In order to teach compressor to extract additional features from $\mathbf{h}$ instead of just preserving its semantics, we apply perturb-and-recover training procedure.
@@
   2. (b)
-    Gaussian noise: after normalising $\mathbf{h}$ with the pre-computed statistics, we inject noise via $\mathbf{h}^{\prime}=\delta\,\mathbf{h}+\sqrt{1-\delta^{2}}\,\varepsilon$ with $\delta=0.7$ and $\varepsilon\sim\mathcal{N}(0,I)$.
-These augmentations force the autoencoder to tolerate partial information loss and promote smooth interpolation between adjacent latents (see Section 5.1).
-Latent-space augmentation.
+    Gaussian noise: after normalising $\mathbf{h}$ with the pre‑computed statistics, we inject noise via $\mathbf{h}^{\prime}=\delta\,\mathbf{h}+\sqrt{1-\delta^{2}}\,\varepsilon$ with $\delta=0.7$ and $\varepsilon\sim\mathcal{N}(0,I)$.
+These augmentations force the autoencoder to tolerate partial information loss and promote smooth interpolation between adjacent latents (see Section˜5.1).
+Latent‑space augmentation.
 We also apply augmentation directly to the latent matrix $\mathbf{z}$.
@@
 Instead, during training we randomly zero out a fixed proportion $p$ of the individual features inside each latent vector.
-This fine-grained sparsification encourages neighbouring features within every latent vector to store redundant cues about one another, so that the representation remains intelligible even when a subset of features is excised, thereby making the manifold more robust to small latent perturbations..
-5 Latent-space properties that facilitate diffusion training
-This section embarks on an exploration of the text-latent manifold, asking which of its intrinsic features govern the performance of a diffusion model.
-All experiments reported here employ autoencoders that compress 128-token texts into 16 latent vectors.
+This fine‑grained sparsification encourages neighbouring features within every latent vector to store redundant cues about one another, so that the representation remains intelligible even when a subset of features is excised, thereby making the manifold more robust to small latent perturbations..
+5 Latent‑space properties that facilitate diffusion training
+This section embarks on an exploration of the text‑latent manifold, asking which of its intrinsic features govern the performance of a diffusion model.
+All experiments reported here employ autoencoders that compress 128‑token texts into 16 latent vectors.
 The autoencoders are trained on the Wikipedia dataset, whereas the diffusion models are optimised on the smaller yet high-quality ROCStories (mostafazadeh2016corpus, ) dataset.
@@
   3. 3.
-    Apply diffusion noise (Equation 1) to get $\mathbf{z}^{\mu}_{t}$ at different time steps $t$ to measure how unseen regions influence diffusion at different noise levels.
+    Apply diffusion noise (Equation˜1) to get $\mathbf{z}^{\mu}_{t}$ at different time steps $t$ to measure how unseen regions influence diffusion at different noise levels.
   4. 4.
@@
   5. 5.
-    Evaluate text plausibility with GPT–2 perplexity (PPL), averaged over 1 00010001\,0001 000 random endpoint pairs.
+    Evaluate text plausibility with GPT–2 perplexity (PPL), averaged over 1 0001\,000 random endpoint pairs.
 Figure 3: PPL of texts decoded from an interpolation of two latents for Cosmos and CE baseline.
-By analysing how perplexity (PPL) evolves with the interpolation coefficient $\mu$ and the diffusion timestep $t$, we obtain a fine-grained picture of manifold smoothness and model robustness far beyond the training support.
-Figure 3 shows that autoencoder trained only with cross-entropy (CE) loss and our robustness-oriented alternative behave identically when $\mu\!\to\!0$ or $\mu\!\to\!1$, confirming that both models successfully reconstruct train-time input latents.
+By analysing how perplexity (PPL) evolves with the interpolation coefficient $\mu$ and the diffusion timestep $t$, we obtain a fine‑grained picture of manifold smoothness and model robustness far beyond the training support.
+Figure 3 shows that autoencoder trained only with cross‑entropy (CE) loss and our robustness‑oriented alternative behave identically when $\mu\!\to\!0$ or $\mu\!\to\!1$, confirming that both models successfully reconstruct train-time input latents.
 As we move away from the endpoints, perplexity rises for both encoders, but it escalates far more rapidly for the CE baseline — clear evidence that the latent manifold learned by our autoencoder is markedly smoother.
@@
 To assess how well the decoder tolerates perturbations, we inject Gaussian noise into latent vectors of real texts: $\mathbf{z}_{\text{noised}}=\mathbf{z}+\sigma\,\boldsymbol{\varepsilon}$, where $\boldsymbol{\varepsilon}\sim\mathcal{N}(0,I)$, decode the perturbed vector to obtain $\mathbf{\hat{w}}$, and measure its BLEU score against the original text $\mathbf{w}$.
-Figure 4 shows that two components contribute most to decoder stability: (i) an explicit MSE loss on the decompressor’s output (§4.2.2), which prevents the final-layer norm from exploding, a common effect observed when training solely with cross-entropy loss (gao2021simcse, ), and (ii) latent masking (§4.2.2), which forces the compressor to distribute information more evenly among latent features and remain resilient to partial dropout.
+Figure 4 shows that two components contribute most to decoder stability: (i) an explicit MSE loss on the decompressor’s output (§4.2.2), which prevents the final‑layer norm from exploding, a common effect observed when training solely with cross-entropy loss (gao2021simcse, ), and (ii) latent masking (§4.2.2), which forces the compressor to distribute information more evenly among latent features and remain resilient to partial dropout.
 Notably, the decoder supplemented with described train-time modifications can reconstruct text almost perfectly even under substantial noise ($\sigma=1$).
@@
 Diffusion robustness during generation.
-Next, we test the diffusion model’s sensitivity to small perturbations introduced mid-trajectory.
+Next, we test the diffusion model’s sensitivity to small perturbations introduced mid‑trajectory.
 For a late generation step $t$, chosen from the second half of the reverse process, where latent representations are already semantically meaningful, we add Gaussian noise with magnitude $\nu$ and continue sampling.
@@
 Table 1: Train–inference mismatch for different autoencoder objectives.
-Configuration MSE$\bigl{(}\hat{\mathbf{z}},E(D(\hat{\mathbf{z}}))\bigr{)}$
-$\mathrm{CE}(w,\hat{w})$ 0.721
 […]
+[…]
+   + MSE(h,h^)($\hat{h}$) 0.619
    + Activation aug. 0.582
@@
 Direct mismatch measurement.
-Table 1 tracks the train–inference mismatch, quantified as the MSE between $\hat{\mathbf{z}}$ and $E\bigl{(}D(\hat{\mathbf{z}})\bigr{)}$.
-Each proposed refinement tightens the gap.
-The consistent drop in error confirms that the enhanced training procedure makes the autoencoder more resilient to diffusion perturbations.
-Together, these results demonstrate that our training recipe simultaneously strengthens the decoder and the diffusion model, thereby shrinking the train–inference gap.
+[…]
 6 Empirical study of latent representations for diffusion modeling
@@
 Evaluation metrics.
-We evaluate the unconditional generation capacity of models with three complementary metrics that, together, capture textual quality, lexical diversity, and distributional fidelity.
-First, we report perplexity (PPL), computed with GPT-2 Large (radford2019language, ).
-Second, we quantify diversity using the score $\mathrm{\textbf{div}}(y)=\prod_{n=2}^{4}\tfrac{\#\,\text{unique }n\text{-gramsin }y}{\#\,n\text{-grams in }y}$, where $y$ denotes the set of generated texts.
-Finally, because low perplexity can be achieved by simple or repetitive texts, we also compute the MAUVE score (pillutla2021mauve, ), which measures the distributional alignment between generated and reference texts and thus offers a broader view of generation quality.
-For every model we generate $1{,}000$ samples and, when calculating MAUVE, draw an equally sized reference subset from the held-out test split.
-The entire evaluation procedure is repeated ten times with different random seeds, and we report the mean and standard deviation of the resulting scores across runs.
+[…]
 Results.
-Table 2 demonstrates that each proposed feature contributes significantly to the model performance.
+Table˜2 demonstrates that each proposed feature contributes significantly to the model performance.
 Remarkably, adding MSE penalty between the predicted and reference BERT activations alone boosts all quality metrics by about 50%.
 Subsequent augmentation of the BERT activations yields a further leap, underscoring the value of intermediate contextual, semantically grounded representations.
-This stage rises MAUVE to 0.767, already achieving the performance obtained when the diffusion model is trained on the full-length, uncompressed BERT representations (0.767 vs.
+This stage rises MAUVE to 0.767, already achieving the performance obtained when the diffusion model is trained on the full‑length, uncompressed BERT representations (0.767 vs.
 0.762 for TEncDM (shabalin2025tencdm, )).
@@
 These results provide a clear answer to the question posed in the introduction of this study: text representations can indeed be mapped into a more compact latent space, where a trained diffusion model performs comparably to traditional token-level counterparts without sacrificing the text generation quality.
-For all subsequent autoencoders we lock in the configuration highlighted in Table 2: the MSE penalty is kept, the random-masking rate is fixed at $0.3$, Gaussian noising is applied with $\delta=0.7$, and latent masking rate is set to $0.4$.
+For all subsequent autoencoders we lock in the configuration highlighted in Table˜2: the MSE penalty is kept, the random‐masking rate is fixed at $0.3$, Gaussian noising is applied with $\delta=0.7$, and latent masking rate is set to $0.4$.
 Additionally, we validate all proposed autoencoder configurations by training the diffusion model on the larger Wikipedia dataset.
-Detailed results of this evaluation are presented in Section C.2.
+Detailed results of this evaluation are presented in LABEL:app:ablation_wikipedia.
 Furthermore, we explore the use of a variational prior in the latent space and observe that it offers no clear advantage.
-A thorough analysis is provided in Section C.1.
+A thorough analysis is provided in Section˜D.1.
 6.2 Impact of scaling the number of latent vectors
-Figure 6: Impact of scaling $N$ on diffusion generation quality.
+Figure 6: Impact of scaling NN on diffusion generation quality.
 In this section, we examine how diffusion generation quality varies when varying the number of latent vectors $N$.
-We keep the training pipeline fixed according to the hyperparameters detailed in Section 6.1.
+We keep the training pipeline fixed according to the hyperparameters detailed in Section˜6.1.
 We do not modify the embedding dimension $d$, because it would necessitate architectural modifications to the diffusion model, making it difficult to isolate the impact of latent space diffusability from changes in model capacity and structure.
-Table 3: Impact of scaling the number of latent vectors $N$ on unconditional generation quality.
-$N$ MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$
+Table 3: Impact of scaling the number of latent vectors NN on unconditional generation quality.
+NN MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$
 Source 0.953 21.7 0.403
@@
 128 0.940.011 26.3.4 0.346.004
-Figure 6 compares baseline cross-entropy compressor to the robustness-oriented compressor introduced in Section 4.2.2.
+Figure˜6 compares baseline cross-entropy compressor to the robustness-oriented compressor introduced in Section˜4.2.2.
 We observe that for CE compressor decrease of $N$ leads to a rapid degradation in quality.
 In contrast, the proposed autoencoder maintains high generation quality even under substantial compression.
-It achieves an 8×8$\times$8 × reduction in latent sequence length while surpassing the quality of uncompressed representations.
+It achieves an $8\times$ reduction in latent sequence length while surpassing the quality of uncompressed representations.
 Table 3 provides further insights by comparing generation quality across different numbers of latent vectors $N$.
-Although configuration $N=16$ matches the quality of BERT representations baseline, we observe a notable quality gap compared to the configuration with $N{=}128$: the MAUVE score decreases from $0.940$ to $0.836$, while perplexity increases from $26.3$ to $30.2$.
-However, adopting a less aggressive compression to $N{=}32$ (a 4×4$\times$4 × compression) results in only minor quality degradation, still significantly outperforming the baseline.
-Based on these observations, together with results from Section 7, we recommend limiting latent compression to approximately 4×4$\times$4 × to optimally balance quality and computational efficiency.
+Although configuration $N=16$ matches the quality of BERT representations baseline, we observe a notable quality gap compared to the configuration with N=128N{=}128: the MAUVE score decreases from $0.940$ to $0.836$, while perplexity increases from $26.3$ to $30.2$.
+However, adopting a less aggressive compression to N=32N{=}32 (a $4\times$ compression) results in only minor quality degradation, still significantly outperforming the baseline.
+Based on these observations, together with results from Section˜7, we recommend limiting latent compression to approximately $4\times$ to optimally balance quality and computational efficiency.
 7 Comparison across generative paradigms
 We evaluate the performance of Cosmos on four distinct text generation tasks: unconditional story generation (ROCStories (mostafazadeh2016corpus, )), text summarization (XSum (narayan2018don, )), detoxification (ParaDetox (logacheva-etal-2022-paradetox, )), and question generation (SQuAD2.0 (minaee2021deep, )).
-For unconditional story generation we employ metrics discussed in Section 6.1.
+For unconditional story generation we employ metrics discussed in Section˜6.1.
 For text summarization and question generation, we report BERTScore (BS) (zhang2019bertscore, ).
@@
 We further assess models’ efficiency by comparing inference time for unconditional generation on sequences with 128 and 512 tokens.
-Additional metrics and implementation details are provided in the Section C.3.
+Additional metrics and implementation details are provided in the Section˜D.2.
 Baselines.
@@
 Autoencoder setup across tasks.
-[…]
+To perform latent-space diffusion, we begin by training autoencoders on Wikipedia dataset, varying the input sequence length.
+For tasks with relatively short input, such as unconditional story generation and text detoxification, we design two variants of the autoencoder that map 128-token input into either $16$ or $128$ latent vectors, resulting in the CosmosN=16N=16 and CosmosN=128N=128 configurations, respectively.
+In contrast, conditional generation tasks such as summarization and question generation require substantially longer input contexts.
+For these settings, we employ a third autoencoder configured to compress 512-token sequences into 128 latent vectors.
+Accordingly, we do not report results for long-context tasks for CosmosN=16N=16, as sych aggressive compression fails to yield plausible outputs in these long-context scenarios.
+We adopt N=128N{=}128 ($4\times$ compression) for long-context tasks as it strikes a favorable balance between generation quality and computational efficiency, as demonstrated in the scaling analysis in Section˜6.2.
 Table 4: Comparison with autoregressive and diffusion baselines across four generative tasks.
@@
 ROCStories XSum ParaDetox SQuAD2.0 Time (s)
-Method MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$ BS $\uparrow$ BLEU $\uparrow$ BS $\uparrow$ $L=128$ $L=512$
-Source text 0.953 21.7 0.403
-GPT2 0.789 20.5 0.252 0.690 0.677 0.680 $\mathbf{1.2_{.1}}$ $42.8_{.1}$
-GPT Neo 0.720 19.9 0.258 0.621 0.610 0.665 $2.6_{.1}$ $45.2_{.2}$
-AR-Diffusion 0.066 41.8 0.101 0.568 0.647 0.569 $226.4_{.2}$ $602_{2.1}$
-DiffuSeq 0.086 50.5 0.124 0.588 0.679 0.563 $215.9_{1.2}$ $1565_{.7}$
-SeqDiffuSeq 0.103 29.3 0.137 0.617 0.688 0.574 $100_{.5}$ $601_{.5}$
-TESS 0.061 22.4 0.163 0.627 0.693 0.667 $1441_{.2}$ $5984_{.0}$
-SEDD 0.598 70.8 0.336 0.576 0.666 0.443 $15.0_{1.1}$ $60.3_{1.0}$
-LD4LG 0.716 30.6 0.331 0.702 0.708 0.641 $27.9_{.5}$ $102_{.2}$
-TEncDM 0.762 29.1 0.295 0.699 0.619 0.703 $29.6_{.1}$ $180_{.3}$
-CosmosN=16 0.836 30.2 0.322 0.654 $5.8_{.1}$
-CosmosN=128 0.940 26.3 0.346 0.704 0.694 0.708 $35.1_{.1}$ $\mathbf{36.6_{.1}}$
+Method MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$ Mem $\downarrow$ BS $\uparrow$ BLEU $\uparrow$ BS $\uparrow$ $L=128$ $L=512$
+Source text 0.953 21.7 0.403 0.365
+GPT2 0.789 20.5 0.252 0.455 0.690 0.677 0.680 $\mathbf{1.2_{.1}}$ $42.8_{.1}$
+GPT Neo 0.720 19.9 0.258 0.469 0.621 0.610 0.665 $2.6_{.1}$ $45.2_{.2}$
+AR-Diffusion 0.066 41.8 0.101 0.540 0.568 0.647 0.569 $226.4_{.2}$ $602_{2.1}$
+DiffuSeq 0.086 50.5 0.124 0.516 0.588 0.679 0.563 $215.9_{1.2}$ $1565_{.7}$
+SeqDiffuSeq 0.103 29.3 0.137 0.663 0.617 0.688 0.574 $100_{.5}$ $601_{.5}$
+TESS 0.061 22.4 0.163 0.550 0.627 0.693 0.667 $1441_{.2}$ $5984_{.0}$
+SEDD 0.598 70.8 0.336 0.325 0.576 0.666 0.443 $15.0_{1.1}$ $60.3_{1.0}$
+LD4LG 0.716 30.6 0.331 0.432 0.702 0.708 0.641 $27.9_{.5}$ $102_{.2}$
+TEncDM 0.762 29.1 0.295 0.438 0.699 0.619 0.703 $29.6_{.1}$ $180_{.3}$
+CosmosN=16 0.836 30.2 0.322 0.394 0.654 $5.8_{.1}$
+CosmosN=128 0.940 26.3 0.346 0.383 0.704 0.694 0.708 $35.1_{.1}$ $\mathbf{36.6_{.1}}$
 Results.
-[…]
-[…]
+As shown in Table˜9, CosmosN=16N=16 achieves strong performance relative to its closest latent diffusion baseline, TEncDM.
+It consistently outperforms TEncDM across most evaluation metrics, with the exception of unconditional perplexity, while offering significantly faster generation.
+Scaling up to a bigger latent representation, CosmosN=128N=128 further advances generation quality.
+On ROCStories, it achieves a MAUVE score of $0.940$, closely approaching the human reference score of $0.953$, and substantially outperforming the best autoregressive baseline (GPT2), which reaches only $0.789$.
+Although CosmosN=128N=128 lags behind GPT2 in perplexity, this gap reflects an evaluation bias: perplexity inherently favors autoregressive models, particularly when assessed using the same decoding objective they were trained on.
+In generation tasks that involve longer input contexts, CosmosN=128N=128 consistently matches or slightly exceeds the performance of both diffusion-based and autoregressive baselines, while substantially reducing inference time.
+This efficiency stems from a core advantage of latent diffusion: unlike autoregressive decoding, where inference time grows linearly with sequence length, diffusion models operate with a fixed number of sampling steps.
+Although this cost is less favorable for short sequences, the benefits become increasingly pronounced as input length grows.
+This favorable scaling behavior makes latent diffusion as a promising direction for future research.
+8 Experiments on OpenWebText
+In this section, we present an empirical evaluation of our proposed method on the large-scale OpenWebText (OWT) (Gokaslan2019OpenWeb, ) dataset.
+8.1 Impact of Diffusion Model Depth
+Table 5: Scaling results on OpenWebText for 128-token generation. Increasing the number of layers in the diffusion model transformer improves generation quality and diversity. Best results in bold.
+Model Size MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$ Mem $\downarrow$
+0.12 B 0.849 97.6 0.492 0.135
+0.25 B 0.914 91.2 0.546 0.124
+0.5 B 0.923 89.7 0.554 0.125
+We train three diffusion models with varying depths of 12, 24, and 48 layers, which correspond to approximately 0.12B, 0.25B, and 0.5B parameters, respectively.
+For these experiments, the autoencoder and latent space configuration are held constant.
+We selected our best-performing autoencoder from preliminary experiments, which utilizes the largest latent space of 128 vectors for 128-token sequences.
+The results presented in Table 5 clearly indicate a positive scaling trend.
+As the number of layers in the diffusion model increases, both generation quality and diversity improve.
+This demonstrates that increasing the capacity of the diffusion component is an effective strategy for enhancing performance, even when the autoencoder’s architecture remains fixed.
+8.2 Comparison on Long-Sequence Generation
+Table 6: Comparison with the GIDD masked diffusion model on OpenWebText for 512-token generation. Our model (CosmosN=512) shows significantly better quality and coherence.
+Model MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$ Mem $\downarrow$
+Source text 0.968 23.2 0.465 0.036
+GIDD 0.286 228.3 0.588 0.112
+TEncDM 0.228 118.6 0.324 0.102
+CosmosN=256 0.263 84.3 0.368 0.156
+CosmosN=512 0.702 55.0 0.319 0.180
+We further assess our model’s capabilities on the challenging task of generating 512-token sequences.
+We compare our model, Cosmos, in two configurations: without compression ($N=512$ latent vectors) and with 2x compression ($N=256$).
+These are benchmarked against TEncDM (shabalin2025tencdm, ), a latent diffusion baseline, and GIDD (von2025generalized, ), a strong discrete diffusion model.
+The results, presented in Table 6, offer several key insights into the advantages of our approach.
+First, our primary model, CosmosN=512, demonstrates a significant performance gap over the discrete diffusion paradigm.
+It outperforms GIDD by a large margin, achieving a MAUVE score that is nearly three times higher (0.792 vs.
+0.286) and a perplexity that is almost four times lower (55.0 vs.
+228.3), indicating superior text quality and coherence.
+Second, the results confirm the expected trade-off between compression and generation fidelity.
+While our compressed model, CosmosN=256, shows a degradation in performance compared to its uncompressed counterpart, it still outperforms TEncDM, a baseline that works with full-length, uncompressed textual representations.
+This result powerfully highlights the efficiency of our latent diffusion framework for text generation.
 9 Conclusion
@@
 Our findings challenge the dominance of token-level models and highlight latent diffusion as a promising direction for building fast, high-quality language models.
+Acknowledgements
+The work was supported by the grant for research centers in the field of AI provided by the Ministry of Economic Development of the Russian Federation in accordance with the agreement  [^00] C313925P4E0002 and the agreement with HSE University № 139-15-2025-009.
+This research was supported in part through computational resources of HPC facilities at HSE University.
 Appendix
-[…]
-section0em2em
 Appendix A Dataset descriptions
@@
 Both context and target sequences are truncated to 40 tokens.
-Appendix B Implementation details
-B.1 Latent diffusion pipeline
+Appendix B Additional Related Work
+In addition to the works discussed in the main paper, a significant body of research has explored diffusion models for text generation. These can be broadly categorized as follows.
+B.1 Gaussian Diffusion on Embeddings
+One line of work focuses on applying Gaussian diffusion directly to word embeddings.
+Models like DiffuSeq [10] and SeqDiffuSeq [45] are designed for sequence-to-sequence text generation tasks.
+DiffuSeq employs a partial noising strategy and is trained end-to-end in a classifier-free manner.
+SeqDiffuSeq utilizes an encoder-decoder Transformer architecture and incorporates self-conditioning and an adaptive noise schedule to enhance performance.
+AR-Diffusion [43] integrates autoregressive principles into the diffusion process to better capture the sequential dependencies in language, allowing tokens on the right to depend on the generated tokens on the left.
+This is achieved by using a dynamic number of denoising steps that varies with the token’s position.
+B.2 Simplex-Based Diffusion
+Another direction explores alternative spaces for the diffusion process.
+TESS [20], for instance, applies a fully non-autoregressive diffusion process on the logit simplex space.
+This approach, combined with a novel self-conditioning technique, demonstrates strong performance on various natural language generation tasks while requiring fewer diffusion steps.
+B.3 Masked Diffusion
+Masked-token diffusion models adopt a BERT-style prediction objective with iterative unmasking, enabling fast and parallel generation.
+For natural language, SEDD [18] established a connection between score-based models and ELBO maximization, demonstrating the effectiveness of the approach.
+Subsequent work, such as simple absorbing-mask diffusion with a clean training recipe, has successfully narrowed the perplexity gap to autoregressive language models and supports flexible semi-autoregressive decoding [34].
+The quality of generation is also significantly boosted by various inference-time techniques.
+These include remasking for iterative correction [41], discrete guidance for improved controllability [35], and hybrid methods that expand self-correction capabilities [40].
+B.4 Latent Diffusion
+In the realm of latent diffusion, several alternative baseline models have been proposed.
+LD4LG [19] learns a continuous diffusion model in the latent space of a pretrained encoder-decoder model.
+This method samples continuous latent representations which are then decoded into natural language, proving effective for both unconditional and conditional text generation.
+Similarly, PLANNER [47] adopts a two-stage recipe where a fine-tuned BERT encoder produces a variational latent code, which is then refined by a diffusion model before being translated into text by a GPT-2 decoder.
+TEncDM [36], on the other hand, operates in the space of pretrained language model encodings, which, unlike embeddings, integrate contextual information.
+Appendix C Implementation details
+C.1 Latent diffusion pipeline
 With a diffusable latent manifold in hand (§4.2.2), we now describe the diffusion model that operates entirely within this space.
@@
 (2): we train a denoiser that receives a noisy latent $\mathbf{z}_{t}$ at diffusion time $t\!\in\![0,1]$ and predicts the corresponding clean latent $\mathbf{z}_{0}$.
-The loss is the simple mean-squared error $\lVert\mathbf{z}_{0}-z_{\theta}(z_{t},t)\rVert_{2}^{2}$, where $z_{\theta}$ shares parameters across all time-steps.
-In line with earlier work [18, 33], we employ self-conditioning: on $50\%$ of training updates the model is fed its own previous estimate of $\mathbf{z}_{0}$ as an additional input.
+The loss is the simple mean‑squared error $\lVert\mathbf{z}_{0}-z_{\theta}(z_{t},t)\rVert_{2}^{2}$, where $z_{\theta}$ shares parameters across all time‑steps.
+In line with earlier work [19, 36], we employ self‑conditioning: on $50\%$ of training updates the model is fed its own previous estimate of $\mathbf{z}_{0}$ as an additional input.
 This iterative refinement enables the model, at inference time, to reuse its past predictions, producing markedly crisper and more coherent text.
@@
 The denoiser itself inherits the TEncDM [36] architecture.
-It is a 12-layer Transformer with 12 attention heads and a hidden size of 768.
+It is a 12‑layer Transformer with 12 attention heads and a hidden size of 768.
 In total, the network comprises roughly $\approx\!130$M parameters.
@@
 For conditional tasks, we initiate the training of our diffusion model by initializing its denoiser weights with those pretrained on an unconditional task.
-B.2 Training and inference configuration
-[…]
+C.2 Training and inference configuration
+All models are trained on 8 NVIDIA A100 GPUs. Detailed training configurations and approximate durations for both CosmosN=128N=128 and CosmosN=16N=16 are provided in Table 7.
 ROCStories Wikipedia XSum SQuAD2.0 Paradetox
@@
 Table 7: Training details for Cosmos across different datasets.
-B.3 Inference time benchmarking configuration
+C.3 Inference time benchmarking configuration
 We measure the generation time for a batch of 512 samples per model on a single NVIDIA A100 GPU using the bfloat16 data type.
@@
 We report the mean and standard deviation across five independent runs.
-Appendix C Additional autoencoder training analysis
-C.1 Marginal utility of variational prior in text autoencoder
+Appendix D Additional autoencoder training analysis
+D.1 Marginal utility of variational prior in text autoencoder
 In this section, we explain why incorporating a variational prior does not significantly enhance the performance of our text autoencoder.
@@
 $\mathcal{L}=\mathrm{CE}(w,\hat{w})\;+\;\mathrm{MSE}(h,\hat{h})\;+\;\beta\,D_{\mathrm{KL}}\!\bigl(\mathcal{N}(z;\mu,\sigma^{2})\,\|\,\mathcal{N}(0,I)\bigr),$
-where $\mathrm{CE}(w,\hat{w})$ is the token-level cross-entropy, $\mathrm{MSE}(h,\hat{h})$ matches reconstructed contextual representations, and $\beta\!\geq\!0$ balances reconstruction fidelity against latent regularisation.
+[…]
 During training, the decoder receives a stochastic latent sample
@@
 Increasing $\beta$ strengthens this pressure, producing a smoother latent geometry at the cost of higher reconstruction error; decreasing $\beta$ does the opposite as encoder tends to push latents away from each other.
-In our experiments, we observed that the performance is highly sensitive to the choice of  $\beta$. Table 6 presents the results for different values of  $\beta$.
+In our experiments, we observed that the performance is highly sensitive to the choice of  $\beta$. Table˜8 presents the results for different values of  $\beta$.
 Table 8: Impact of the KL weight $\beta$ on diffusion generation quality. Arrows indicate the preferred direction.
@@
 This limited improvement can be attributed to the decoder’s robustness to Gaussian noise in the latent space.
-As illustrated in Figure 4, the decoder maintains reasonable performance even when sampling from the prior with $\sigma=1$, indicating that the latent space is sufficiently robust without strong KL regularization.
+As illustrated in Figure˜4, the decoder maintains reasonable performance even when sampling from the prior with $\sigma=1$, indicating that the latent space is sufficiently robust without strong KL regularization.
 Alternative regularisers, notably latent masking, yield larger and more robust improvements.
 We therefore omit the KL term in the final model and rely solely on latent masking to shape the representation space.
-C.2 Empirical analysis of autoencoder training regimes on Wikipedia dataset
-Table 7: Comparison of autoencoder training regimes on text generation quality on Wikipedia. Features are added cumulatively from top to bottom.
-Configuration MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$
-CE($\bf{w,\hat{w}}$) $0.056$ $385.6$ $0.590$
-+ MSE($\bf{h,\hat{h}}$) $0.069$ $344.4$ $0.625$
-+ Random masking (rate)
-   0.3 $0.081$ $289.1$ $0.594$
-+ Gaussian noising ($\delta$)
-   0.7 $0.103$ $205.0$ $0.588$
-+ Latent dropout (rate)
-   0.4 $0.112$ $186.7$ $0.581$
-We also conduct a large-scale ablation on the Wikipedia dataset with $N=16$ latents and max sequence length $L=128$, following a setup analogous to that in section 6.1.
-Results, presented in Table 7, reveal trends consistent with those observed in the small-scale ablation.
-Notably, applying the proposed training modifications leads to substantial improvements in evaluation metrics.
-The final model approaches the performance of the TEncDM baseline [33], while operating in a latent space that is 8×8$\times$8 × smaller.
-Additionally, we investigate how scaling the latent space length on Wikipedia affects performance.
-We experiment with latent sequence lengths ranging from $16$ to $128$.
-As expected, we observe steady performance improvements with longer latent sequences.
-Results are provided in Table 8.
-Table 8: Impact of scaling the number of latent vectors $N$ on unconditional generation quality on Wikipedia.
-$N$ MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$
-Source 0.953 21.7 0.403
-BERT repr. 0.109 173.1 0.562
-16 0.112 186.7 0.581
-32 0.149 134.4 0.579
-64 0.157 128.1 0.582
-128 0.173 118.2 0.592
-C.3 Additional details of comparison across generative paradigms
+D.2 Additional details of comparison across generative paradigms
 Table 9: Comparison with autoregressive and diffusion baselines across three generative tasks. The best-performing scores are shown in bold, while the second-best scores are underlined.
@@
 For ParaDetox, we additionally include the J-score, which is defined as the product of style accuracy, fluency, and content preservation.
-Extended results are summarized in Table 9.
-Appendix D Limitations
+Extended results are summarized in Table˜9.
+Appendix E Limitations
 There are several limitations to our study that point to promising directions for future work.
@@
 Finally, we report results with relatively small backbones of roughly 130M parameters to keep ablations rapid and compute budgets fair; scaling the architecture is an orthogonal engineering effort that is expected to reinforce the empirical trends observed here.
-Appendix E Societal Impact
+Appendix F Societal Impact
 Our research introduces an alternative modeling approach for language, centered on latent continuous diffusion.
 We hold the view that this method does not bring about significant new societal risks exceeding those already connected with current language models.
-Appendix F Generation examples
+Appendix G Generation examples
 To give a qualitative sense of model behaviour, we first present representative unconditional generations: Table 10 shows random samples from CosmosN=128 and the TEncDM baseline on ROCStories, while Table 11 does the same for Wikipedia.

### v3

@@
-Compressed and Smooth Latent Space for Text Diffusion Modeling
+Cosmos: Compressed and Smooth Latent Space for Text Diffusion Modeling
 Abstract
@@
 1 Introduction
-Autoregressive (AR) language models are a de facto gold standard for text generation (radford2018improving, ; radford2019language, ; brown2020gpt3, ).
-By factorizing the probability of a sequence into a product of conditional token probabilities, they transform the global generation task into a series of local next-token prediction tasks that can be optimized efficiently via teacher forcing (williams1989learning, ).
+Autoregressive (AR) language models are a de facto gold standard for text generation (radford2018improving; radford2019language; brown2020gpt3).
+By factorizing the probability of a sequence into a product of conditional token probabilities, they transform the global generation task into a series of local next-token prediction tasks that can be optimized efficiently via teacher forcing (williams1989learning).
 The same sequential factorization, however, creates several structural bottlenecks.
 First, decoding is inherently sequential: the time to produce a sequence grows linearly with its length because each token must await the completion of its predecessor.
-Second, the process suffers from exposure bias: a single early mistake contaminates the context for all subsequent predictions and cannot be corrected without restarting generation (ranzato2015sequence, ; bengio2015scheduled, ; he2019exposure, ).
-Third, maximizing local log-likelihood encourages the model to privilege fluency over factual accuracy, resulting in the well-documented phenomenon of hallucination (maynez2020faithfulness, ; ji2023survey, ).
-Finally, because decisions are made token-by-token, the model lacks an explicit global plan and therefore struggles to maintain long-range logical or narrative coherence (ye2024beyond, ).
-In stark contrast, computer vision has been reshaped by diffusion models (ho2020ddpm, ; nichol2021improved, ), particularly by latent diffusion (rombach2022ldm, ; blattmann2023stable, ; andreas2023, ), which first compresses a high-resolution image or video into a compact latent representation and then performs the diffusion process in that space.
+Second, the process suffers from exposure bias: a single early mistake contaminates the context for all subsequent predictions and cannot be corrected without restarting generation (ranzato2015sequence; bengio2015scheduled; he2019exposure).
+Third, maximizing local log-likelihood encourages the model to privilege fluency over factual accuracy, resulting in the well-documented phenomenon of hallucination (maynez2020faithfulness; ji2023survey).
+Finally, because decisions are made token-by-token, the model lacks an explicit global plan and therefore struggles to maintain long-range logical or narrative coherence (ye2024beyond).
+In stark contrast, computer vision has been reshaped by diffusion models (ho2020ddpm; nichol2021improved), particularly by latent diffusion (rombach2022ldm; blattmann2023stable; andreas2023), which first compresses a high-resolution image or video into a compact latent representation and then performs the diffusion process in that space.
 Operating in a low-dimensional latent manifold reduces computational cost by orders of magnitude and enables breathtaking advances in image and video synthesis.
-Yet recent works (kouzelis2025eq, ; skorokhodov2025improving, ) also show that diffusion is highly sensitive to the geometry of the latent space: poorly designed latent representations can destabilize training and degrade sample quality, underscoring the need for principled mechanisms to construct diffusable [^1] Diffusable representations refer to latent spaces that permit effective modeling by a diffusion process.
+Yet recent works (kouzelis2025eq; skorokhodov2025improving) also show that diffusion is highly sensitive to the geometry of the latent space: poorly designed latent representations can destabilize training and degrade sample quality, underscoring the need for principled mechanisms to construct diffusable [^1] Diffusable representations refer to latent spaces that permit effective modeling by a diffusion process.
 representations.
@@
 Early efforts to bring latent diffusion models into natural language generation treat the diffusion module mainly as a pre-processor that supplies conditioning vectors for an autoregressive decoder.
-LD4LG (lovelace2023latent, ) trains a diffusion model on compressed hidden states from BART, then feeds the generated latent vectors into the BART decoder to produce text.
-PLANNER (zhang2023planner, ) adopts a similar two-stage recipe: a fine-tuned BERT encoder produces a 16-token variational latent code, a diffusion model refines that code, and a GPT-2 decoder generates text from this code.
+LD4LG (lovelace2023latent) trains a diffusion model on compressed hidden states from BART, then feeds the generated latent vectors into the BART decoder to produce text.
+PLANNER (zhang2023planner) adopts a similar two-stage recipe: a fine-tuned BERT encoder produces a 16-token variational latent code, a diffusion model refines that code, and a GPT-2 decoder generates text from this code.
 While both systems improve controllability, their heavy reliance on powerful autoregressive decoders makes it difficult to isolate and evaluate the intrinsic generative capacity of the latent-space diffusion itself.
 A complementary line of work applies diffusion directly to the continuous embeddings of pretrained encoders.
-TEncDM (shabalin2025tencdm, ) demonstrates that full-length BERT representations can be modeled with Gaussian diffusion and decoded into coherent text without an intermediate autoregressive step.
+TEncDM (shabalin2025tencdm) demonstrates that full-length BERT representations can be modeled with Gaussian diffusion and decoded into coherent text without an intermediate autoregressive step.
 Crucially, however, TEncDM retains the original sequence dimensionality, leaving open a question: How aggressively can such representations be compressed before generation quality collapses?
@@
 3 Preliminary
-Diffusion models (ho2020ddpm, ; nichol2021improved, ) learn a data distribution by reversing a progressive noising process.
+Diffusion models (ho2020ddpm; nichol2021improved) learn a data distribution by reversing a progressive noising process.
 Given a clean sample $z_{0}\sim p_{\text{data}}$, the forward dynamic corrupts the input with Gaussian noise whose magnitude is controlled by a continuous time index $t\in[0,1]$:
@@
 3.1 Latent diffusion
-Latent Diffusion Models (LDMs) (rombach2022ldm, ; andreas2023, ) improve both compute efficiency and quality by learning the diffusion process in a compact latent space rather than in pixel or token space.
+Latent Diffusion Models (LDMs) (rombach2022ldm; andreas2023) improve both compute efficiency and quality by learning the diffusion process in a compact latent space rather than in pixel or token space.
 An autoencoder with encoder $E$ and decoder $D$ is first trained to achieve high-fidelity reconstruction, $\mathbf{\hat{w}}~=~D\!\bigl(E(\mathbf{w})\bigr)~\approx~\mathbf{w},$ where $\mathbf{z}=E(\mathbf{w})$ denotes the low-dimensional latent.
@@
 Frozen text encoder.
-We initialise the pipeline with a pretrained contextual encoder $E_{\text{text}}$ (BERT-base (devlin2019bert, ) by default) that remains frozen during all subsequent training stages.
+We initialise the pipeline with a pretrained contextual encoder $E_{\text{text}}$ (BERT-base (devlin2019bert) by default) that remains frozen during all subsequent training stages.
 For an input sequence of $L$ tokens $\mathbf{w}=(w_{1},\dots,w_{L})$, the encoder produces a matrix of hidden states $\mathbf{h}=E_{\text{text}}(\mathbf{w})\in~\mathbb{R}^{L\times d}$, where $d=768$.
@@
 All experiments reported here employ autoencoders that compress 128‑token texts into 16 latent vectors.
-The autoencoders are trained on the Wikipedia dataset, whereas the diffusion models are optimised on the smaller yet high-quality ROCStories (mostafazadeh2016corpus, ) dataset.
+The autoencoders are trained on the Wikipedia dataset, whereas the diffusion models are optimised on the smaller yet high-quality ROCStories (mostafazadeh2016corpus) dataset.
 Across all experiments, we generate 1 000 samples and report each metric as the average over these samples.
@@
 Put differently, the degree to which a locally estimated score function extends to unexplored territory is dictated by manifold smoothness.
-Inspired by PLANNER (zhang2023planner, ), we investigate this property by conducting the following experiment, which simultaneously evaluates performance at points the model does not encounter during training.
+Inspired by PLANNER (zhang2023planner), we investigate this property by conducting the following experiment, which simultaneously evaluates performance at points the model does not encounter during training.
   1. 1.
@@
 To assess how well the decoder tolerates perturbations, we inject Gaussian noise into latent vectors of real texts: $\mathbf{z}_{\text{noised}}=\mathbf{z}+\sigma\,\boldsymbol{\varepsilon}$, where $\boldsymbol{\varepsilon}\sim\mathcal{N}(0,I)$, decode the perturbed vector to obtain $\mathbf{\hat{w}}$, and measure its BLEU score against the original text $\mathbf{w}$.
-Figure 4 shows that two components contribute most to decoder stability: (i) an explicit MSE loss on the decompressor’s output (§4.2.2), which prevents the final‑layer norm from exploding, a common effect observed when training solely with cross-entropy loss (gao2021simcse, ), and (ii) latent masking (§4.2.2), which forces the compressor to distribute information more evenly among latent features and remain resilient to partial dropout.
+Figure 4 shows that two components contribute most to decoder stability: (i) an explicit MSE loss on the decompressor’s output (§4.2.2), which prevents the final‑layer norm from exploding, a common effect observed when training solely with cross-entropy loss (gao2021simcse), and (ii) latent masking (§4.2.2), which forces the compressor to distribute information more evenly among latent features and remain resilient to partial dropout.
 Notably, the decoder supplemented with described train-time modifications can reconstruct text almost perfectly even under substantial noise ($\sigma=1$).
@@
 For the ablation study, we compress text from 128 BERT representations down to 16 latent vectors.
-All autoencoders are trained on the Wikipedia dataset, while the diffusion models are trained on the ROCStories (mostafazadeh2016corpus, ) dataset.
+All autoencoders are trained on the Wikipedia dataset, while the diffusion models are trained on the ROCStories (mostafazadeh2016corpus) dataset.
 Table 2: Comparison of autoencoder training regimes on text generation quality.
@@
 This stage rises MAUVE to 0.767, already achieving the performance obtained when the diffusion model is trained on the full‑length, uncompressed BERT representations (0.767 vs.
-0.762 for TEncDM (shabalin2025tencdm, )).
+0.762 for TEncDM (shabalin2025tencdm)).
 Finally, dropping random features from latent vectors pushes MAUVE beyond 0.83, lowers perplexity to 30.2, keeping diversity essentially unchanged.
@@
 For all subsequent autoencoders we lock in the configuration highlighted in Table˜2: the MSE penalty is kept, the random‐masking rate is fixed at $0.3$, Gaussian noising is applied with $\delta=0.7$, and latent masking rate is set to $0.4$.
-Additionally, we validate all proposed autoencoder configurations by training the diffusion model on the larger Wikipedia dataset.
-Detailed results of this evaluation are presented in LABEL:app:ablation_wikipedia.
-Furthermore, we explore the use of a variational prior in the latent space and observe that it offers no clear advantage.
-A thorough analysis is provided in Section˜D.1.
+Additionally, we explore the use of a variational prior in the latent space and observe that it offers no clear advantage. A thorough analysis is provided in Section˜D.1.
 6.2 Impact of scaling the number of latent vectors
@@
 7 Comparison across generative paradigms
-We evaluate the performance of Cosmos on four distinct text generation tasks: unconditional story generation (ROCStories (mostafazadeh2016corpus, )), text summarization (XSum (narayan2018don, )), detoxification (ParaDetox (logacheva-etal-2022-paradetox, )), and question generation (SQuAD2.0 (minaee2021deep, )).
+We evaluate the performance of Cosmos on four distinct text generation tasks: unconditional story generation (ROCStories (mostafazadeh2016corpus)), text summarization (XSum (narayan2018don)), detoxification (ParaDetox (logacheva-etal-2022-paradetox)), and question generation (SQuAD2.0 (minaee2021deep)).
 For unconditional story generation we employ metrics discussed in Section˜6.1.
-For text summarization and question generation, we report BERTScore (BS) (zhang2019bertscore, ).
+For text summarization and question generation, we report BERTScore (BS) (zhang2019bertscore).
 The detoxification task is evaluated using BLEU-4 (BLEU).
@@
 We benchmark our proposed method against a diverse set of generative paradigms.
-These include Gaussian diffusion on embeddings (DiffuSeq (gong2022diffuseq, ), SeqDiffuSeq (yuan2022seqdiffuseq, ), AR-Diffusion (wu2023ar, )), simplex-based diffusion (TESS (mahabadi2023tess, )), masked diffusion (SEDD (lou2023discrete, )), alternative latent diffusion baselines (LD4LG (lovelace2023latent, ), TEncDM (shabalin2025tencdm, )), and autoregressive models (GPT-2 (radford2019language, ) and GPT-Neo (gpt-neo, )).
+These include Gaussian diffusion on embeddings (DiffuSeq (gong2022diffuseq), SeqDiffuSeq (yuan2022seqdiffuseq), AR-Diffusion (wu2023ar)), simplex-based diffusion (TESS (mahabadi2023tess)), masked diffusion (SEDD (lou2023discrete)), alternative latent diffusion baselines (LD4LG (lovelace2023latent), TEncDM (shabalin2025tencdm)), and autoregressive models (GPT-2 (radford2019language) and GPT-Neo (gpt-neo)).
 All models operate within a comparable parameter scale, ranging from approximately 100M to 200M.
@@
 8 Experiments on OpenWebText
-In this section, we present an empirical evaluation of our proposed method on the large-scale OpenWebText (OWT) (Gokaslan2019OpenWeb, ) dataset.
+In this section, we present an empirical evaluation of our proposed method on the large-scale OpenWebText (OWT) (Gokaslan2019OpenWeb) dataset.
 8.1 Impact of Diffusion Model Depth
@@
 We compare our model, Cosmos, in two configurations: without compression ($N=512$ latent vectors) and with 2x compression ($N=256$).
-These are benchmarked against TEncDM (shabalin2025tencdm, ), a latent diffusion baseline, and GIDD (von2025generalized, ), a strong discrete diffusion model.
+These are benchmarked against TEncDM (shabalin2025tencdm), a latent diffusion baseline, and GIDD (von2025generalized), a strong discrete diffusion model.
 The results, presented in Table 6, offer several key insights into the advantages of our approach.
@@
 ROCStories
-This dataset [24] consists of five-sentence stories and serves as a well-established, small-scale, unconditional benchmark in language diffusion research.
+This dataset [rocstories] consists of five-sentence stories and serves as a well-established, small-scale, unconditional benchmark in language diffusion research.
 It contains a total of 98,161 instances, of which 88,161 are used for training, 10,000 for validation.
 Wikipedia
-For large-scale experiments, we use the English Wikipedia subset from the ROOTS corpus [15]. The resulting dataset comprises 8.6 million sequences, with 50,000 held out for validation.
+For large-scale experiments, we use the English Wikipedia subset from the ROOTS corpus [laurenccon2022bigscience].
+The resulting dataset comprises 8.6 million sequences, with 50,000 held out for validation.
 XSum
-This dataset [26] is used for abstractive summarization and comprises 204,000 BBC news articles paired with summaries covering diverse topics (e.g., sports, politics).
+This dataset [xsum-emnlp] is used for abstractive summarization and comprises 204,000 BBC news articles paired with summaries covering diverse topics (e.g., sports, politics).
 It includes 204,045 training instances, 11,332 validation instances, and 11,334 test instances.
@@
 SQuAD2.0
-This question-answering dataset [31] requires identifying the text span that answers a given question or indicating that no answer is possible from the provided context.
+This question-answering dataset [rajpurkar-etal-2018-know] requires identifying the text span that answers a given question or indicating that no answer is possible from the provided context.
 For generative modeling, we reverse the task: given a context and an answer, the model generates the corresponding question.
@@
 ParaDetox
-For small-scale conditional generation experiments, we use the ParaDetox dataset [17], which comprises 19,766 pairs of toxic and neutral comments.
+For small-scale conditional generation experiments, we use the ParaDetox dataset [logacheva-etal-2022-paradetox], which comprises 19,766 pairs of toxic and neutral comments.
 Both context and target sequences are truncated to 40 tokens.
@@
 One line of work focuses on applying Gaussian diffusion directly to word embeddings.
-Models like DiffuSeq [10] and SeqDiffuSeq [45] are designed for sequence-to-sequence text generation tasks.
+Models like DiffuSeq [gong2022diffuseq] and SeqDiffuSeq [yuan2022seqdiffuseq] are designed for sequence-to-sequence text generation tasks.
 DiffuSeq employs a partial noising strategy and is trained end-to-end in a classifier-free manner.
 SeqDiffuSeq utilizes an encoder-decoder Transformer architecture and incorporates self-conditioning and an adaptive noise schedule to enhance performance.
-AR-Diffusion [43] integrates autoregressive principles into the diffusion process to better capture the sequential dependencies in language, allowing tokens on the right to depend on the generated tokens on the left.
+AR-Diffusion [wu2023ar] integrates autoregressive principles into the diffusion process to better capture the sequential dependencies in language, allowing tokens on the right to depend on the generated tokens on the left.
 This is achieved by using a dynamic number of denoising steps that varies with the token’s position.
@@
 Another direction explores alternative spaces for the diffusion process.
-TESS [20], for instance, applies a fully non-autoregressive diffusion process on the logit simplex space.
+TESS [mahabadi2023tess], for instance, applies a fully non-autoregressive diffusion process on the logit simplex space.
 This approach, combined with a novel self-conditioning technique, demonstrates strong performance on various natural language generation tasks while requiring fewer diffusion steps.
@@
 Masked-token diffusion models adopt a BERT-style prediction objective with iterative unmasking, enabling fast and parallel generation.
-For natural language, SEDD [18] established a connection between score-based models and ELBO maximization, demonstrating the effectiveness of the approach.
-Subsequent work, such as simple absorbing-mask diffusion with a clean training recipe, has successfully narrowed the perplexity gap to autoregressive language models and supports flexible semi-autoregressive decoding [34].
+For natural language, SEDD [lou2023discrete] established a connection between score-based models and ELBO maximization, demonstrating the effectiveness of the approach.
+Subsequent work, such as simple absorbing-mask diffusion with a clean training recipe, has successfully narrowed the perplexity gap to autoregressive language models and supports flexible semi-autoregressive decoding [sahoo2024simple].
 The quality of generation is also significantly boosted by various inference-time techniques.
-These include remasking for iterative correction [41], discrete guidance for improved controllability [35], and hybrid methods that expand self-correction capabilities [40].
+These include remasking for iterative correction [wang2025remasking], discrete guidance for improved controllability [schiff2024simple], and hybrid methods that expand self-correction capabilities [von2025generalized].
 B.4 Latent Diffusion
 In the realm of latent diffusion, several alternative baseline models have been proposed.
-LD4LG [19] learns a continuous diffusion model in the latent space of a pretrained encoder-decoder model.
+LD4LG [lovelace2023latent] learns a continuous diffusion model in the latent space of a pretrained encoder-decoder model.
 This method samples continuous latent representations which are then decoded into natural language, proving effective for both unconditional and conditional text generation.
-Similarly, PLANNER [47] adopts a two-stage recipe where a fine-tuned BERT encoder produces a variational latent code, which is then refined by a diffusion model before being translated into text by a GPT-2 decoder.
-TEncDM [36], on the other hand, operates in the space of pretrained language model encodings, which, unlike embeddings, integrate contextual information.
+Similarly, PLANNER [zhang2023planner] adopts a two-stage recipe where a fine-tuned BERT encoder produces a variational latent code, which is then refined by a diffusion model before being translated into text by a GPT-2 decoder.
+TEncDM [shabalin2025tencdm], on the other hand, operates in the space of pretrained language model encodings, which, unlike embeddings, integrate contextual information.
 Appendix C Implementation details
@@
 The loss is the simple mean‑squared error $\lVert\mathbf{z}_{0}-z_{\theta}(z_{t},t)\rVert_{2}^{2}$, where $z_{\theta}$ shares parameters across all time‑steps.
-In line with earlier work [19, 36], we employ self‑conditioning: on $50\%$ of training updates the model is fed its own previous estimate of $\mathbf{z}_{0}$ as an additional input.
+In line with earlier work [lovelace2023latent, shabalin2025tencdm], we employ self‑conditioning: on $50\%$ of training updates the model is fed its own previous estimate of $\mathbf{z}_{0}$ as an additional input.
 This iterative refinement enables the model, at inference time, to reuse its past predictions, producing markedly crisper and more coherent text.
-We adopt the noise schedule introduced by TEncDM [36]: $\alpha_{t}=\frac{1}{1+\tan(t\pi/2)^{2}\cdot d^{2}}$, where parameter $d$ controls he rate at which noise is injected during the diffusion steps.
+We adopt the noise schedule introduced by TEncDM [shabalin2025tencdm]: $\alpha_{t}=\frac{1}{1+\tan(t\pi/2)^{2}\cdot d^{2}}$, where parameter $d$ controls he rate at which noise is injected during the diffusion steps.
 At inference time we use Euler solver with 200 uniform steps.
-The denoiser itself inherits the TEncDM [36] architecture.
+The denoiser itself inherits the TEncDM [shabalin2025tencdm] architecture.
 It is a 12‑layer Transformer with 12 attention heads and a hidden size of 768.
@@
 We also present extended results for our conditional generation tasks using a broader variety of evaluation metrics.
-For XSum and SQuAD2.0, we report ROUGE scores [16], a standard metric that evaluates text quality based on n-gram overlap with reference outputs.
+For XSum and SQuAD2.0, we report ROUGE scores [lin2004rouge], a standard metric that evaluates text quality based on n-gram overlap with reference outputs.
 For ParaDetox, we additionally include the J-score, which is defined as the product of style accuracy, fluency, and content preservation.
@@
 What we witnessed here today was a miracle, and I want you to acknowledge that. what we witnessed here today was a a miracle, and i want to acknowledge that.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: We claim that CoSmoS learns a compressed and smooth latent representation through carefully designed autoencoder objectives, enabling high-quality text generation with reduced inference time.
+Empirically, CoSmoS matches or surpasses traditional token-level diffusion and autoregressive baselines across a diverse set of generative tasks.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: The discussion of limitations is located in the Appendix˜E.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: We do not have theorems in the paper.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: We thoroughly describe our method in Section˜4.2.2 and provide a complete set of hyperparameters in Appendix˜C.
+We use publicly available datasets and will make our code publicly available upon publication.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We use publicly available datasets. We add code to the supplementary materials and will make our code publicly available upon publication.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: All details could be found in Appendix˜A and Appendix˜C.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34. Justification: Following prior work, we do not report error bars for comparison results across all tasks.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: These details could be found in Appendix˜C.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: Our research does not involve human participants and does not introduce any new datasets; all data used are publicly available.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We discuss societal impact in Appendix˜F.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [No]
+  54. Justification: We do not release any pretrained models or datasets.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We provide all the information about the datasets in Appendix˜A. We also cite the BERT paper whose representations we use.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: We release only our source code which we thoroughly comment.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: Our study does not rely on crowdsourced data and does not involve any research with human participants.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: Our study does not involve crowdsourcing or any research involving human participants.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: LLMs are not utilized in any of the scenarios described.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.